### v1

# Revisiting Robust RAG: Do We Still Need Complex Robust Training in the Era of Powerful LLMs?

Abstract
Retrieval-augmented generation (RAG) systems often suffer from performance degradation when encountering noisy or irrelevant documents, driving researchers to develop sophisticated training strategies to enhance their robustness against such retrieval noise. However, as large language models (LLMs) continue to advance, the necessity of these complex training methods is increasingly questioned. In this paper, we systematically investigate whether complex robust training strategies remain necessary as model capacity grows. Through comprehensive experiments spanning multiple model architectures and parameter scales, we evaluate various document selection methods and adversarial training techniques across diverse datasets. Our extensive experiments consistently demonstrate that as models become more powerful, the performance gains brought by complex robust training methods drop off dramatically. We delve into the rationale and find that more powerful models inherently exhibit superior confidence calibration, better generalization across datasets (even when trained with randomly selected documents), and optimal attention mechanisms learned with simpler strategies. Our findings suggest that RAG systems can benefit from simpler architectures and training strategies as models become more powerful, enabling more scalable applications with minimal complexity.
Revisiting Robust RAG: Do We Still Need Complex Robust Training
in the Era of Powerful LLMs?
Hanxing Ding1,2thanks:   Equal contributions. Shuchang Tao11footnotemark: 1  Liang Pang1thanks:   Corresponding author Zihao Wei1,2 Liwei Chen3Kun XuHuawei Shen1,2 Xueqi Cheng1,2 1Key Laboratory of AI Safety, Institute of Computing Technology, Chinese Academy of Sciences 2 University of Chinese Academy of Sciences  3 Kuaishou Technology {dinghanxing18s, pangliang, weizihao22z, shenhuawei, cxq}@ict.ac.cn
"Entities should not be multiplied unnecessarily."
— Occam’s razor
1 Introduction
Modern LLMs excel in various NLP tasks, such as text generation, knowledge reasoning, and question answering Ye et al. (2024); Tao et al. (2024); Zhao et al. (2023). Their contextual relevance and coherence make them valuable tools across domains. Despite their broad applications, LLMs struggle with external knowledge integration, particularly in RAG systems where poor retrieval results can lead to inaccurate or misleading responses Xu et al. (2024); Tan et al. (2024); Ding et al. (2024). This issue exposes a key limitation in the robustness of LLMs with imperfect retrieval results.
Figure 1: Comparison of EM scores of different document selection strategies on the TriviaQA dataset. As the model capabilities increase, the performance gap ($\Delta$) between the best and worst strategies among the four training methods decreases from 14.65% to 4.36%.
To address this challenge, researchers have proposed various robust training methods to enhance RAG’s resilience against noisy contexts. Approaches mainly focus on selecting high-quality documents or incorporating additional adversarial loss regularization terms during training Yoran et al. (2024); Fang et al. (2024); Krueger et al. (2021); Zhu et al. (2024a); Jin et al. (2024); Zeng et al. (2024). For instance, RetRobust Yoran et al. (2024) suggests using a strategic mixture of relevant and irrelevant documents to train LLMs to withstand noisy contexts. RAAT Fang et al. (2024) and ATM Zhu et al. (2024a) introduce adversarial regularization terms to encourage consistent model performance across different document contexts, making the model less sensitive to retrieval noise.
Given the increasing capabilities of LLMs, a fundamental question emerges: do sophisticated robust training strategies remain necessary as model capacity grows? We first conduct a simple preliminary experiment to investigate this inquiry. As shown in Figure 1, we compare the EM scores of four different training strategies on the TriviaQA dataset. Notably, the performance gap ($\Delta$) between the best and worst strategies shrinks dramatically from 14.65% to 4.36% as the model capacity increases. This significant diminishing gap suggests that powerful LLMs might naturally possess enhanced robustness, potentially challenging the necessity of complex robust training techniques.
In this paper, we present a systematic investigation into the necessity of sophisticated robust training strategies as model capabilities advance. Through extensive experiments across various language models with different capabilities levels, we uncover a counterintuitive phenomenon: while models with limited capabilities benefit from high-quality document selection and complex adversarial losses, these benefits diminish substantially as model capability increases. Notably, for highly capable models, training with randomly selected or even seemingly irrelevant documents can achieve comparable or superior performance to sophisticated training strategies.
We further conduct an in-depth study on this counter-intuitive phenomenon through a systematic analysis. Through inherent confidence calibration analysis, we discover that newer powerful models naturally exhibit superior discrimination between correct and incorrect responses. Moving to external performance, we find these models demonstrate remarkable cross-dataset generalization with basic random document training. Our visualization of attention patterns further reveals the underlying mechanism: powerful models inherently learn effective attention allocation even with simple training. These findings, from internal capabilities to external performance to mechanism analysis, indicate that powerful models may naturally possess fundamental desired capabilities, potentially reducing the need for complex training strategies.
We provide insights for future RAG development, suggesting: simplified architecture design principles for powerful models, opportunities for scalable open-domain applications with minimal supervision, and theoretical perspectives on how training requirements evolve with model scaling.
The main contributions of this work are:
  • Finding: We reveal that as model capabilities increase, the benefits of sophisticated training strategies diminish substantially.
  • Rationale: Through systematic analysis, we trace this finding to powerful models’ inherent capabilities, demonstrating that they naturally possess strong confidence calibration, generalization, and effective attention mechanisms, regardless of training sophistication.
  • Insights: We provide practical guidelines for future RAG development, advocating for architectural simplification with powerful models and highlighting new theoretical perspectives on model scaling laws.
2 Background
2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) enhances language model performance by incorporating external knowledge retrieval into the generation process. Given a query $q$, a retriever $\mathcal{R}$ selects a set of $k$ relevant document sets $\mathcal{D}=\{d_{1},d_{2},\dots,d_{k}\}$ from a large corpus, providing additional factual grounding for the model. These retrieved documents serve as context to improve the model’s reasoning and factual accuracy. The language model $G$ then generates an answer $a$ based on the augmented input, following the conditional probability formulation:
$G(a\mid q,\mathcal{D})=\prod_{t=1}^{T}P(a_{t}\mid a_{<t},q,\mathcal{D};\theta),$ (1)
where $a=(a_{1},a_{2},\dots,a_{T})$ represents the generated sequence, and $\theta$ denotes the model parameters. The retrieval-augmented setup enables the model to dynamically condition its responses on external information, improving its ability to generate factually consistent answers. However, RAG models remain susceptible to retrieval noise, such as irrelevant or misleading documents, which can degrade performance and lead to hallucinations.
2.2 Robust Training for RAG
Document Selection Strategy
A common strategy for robust training in RAG models is to select high-quality documents from either the golden set or top-ranked passages for fine-tuning. Traditionally, supervised fine-tuning (SFT) relies on high-quality golden data, where each query is paired with a carefully curated document containing the correct answer. This ensures that the model learns from reliable sources, leading to improved factual accuracy. Naturally, leveraging such golden documents in RAG training is a straightforward and effective approach.
However, in real-world scenarios, the golden document may not always be retrieved due to limitations in the retriever. In such cases, training only on golden passages lead to unrealistic expectations during inference. A more practical approach is to incorporate top-1 retrieved documents, which better reflect what the model will encounter at test time. Even if these documents are imperfect, training on them enhances the model’s adaptability. Further, drawing from a mixture of top-ranked, lower-ranked, or randomly sampled passages, as suggested by RetRobust Yoran et al. (2024), exposes the model to diverse retrieval conditions, improving its robustness to noisy or suboptimal results. This strategy ensures that RAG models generalize better across different retrieval settings, making them more reliable in real-world applications.
Adversarial Loss Design
To enhance the robustness of RAG models against retrieval variations, robust training for RAG introduces a general training framework that combines standard SFT with a robustness-enforcing regularization term. During training, the model learns across multiple retrieval environments, which may include golden documents, top-k retrieved results, or adversarially perturbed documents. The overall training objective consists of two components: the first part is the standard training loss across different retrieval environments, and the second part is a regularization term designed to reduce model performance sensitivity to retrieval variations. Different forms of regularization strategies are detailed in Appendix B.
3 Experimental Setups
In this section, we introduce the dataset and evaluation metrics we used to assess the robustness performance of various robust training strategies.
3.1 Datasets and Evaluation Metrics
For our experiment, we evaluate on four widely-used question answering datasets: (1) single-hop QA, including NaturalQuestions (NQ) Kwiatkowski et al. (2019) and WebQuestions Berant et al. (2013); and (2) multi-hop QA, including TriviaQA Joshi et al. (2017) and HotpotQA Yang et al. (2018). All experimental results are evaluated on their development splits using the Exact Match (EM) and F1 metrics. Detailed statistics of these datasets are listed in Appendix C.1 Table 2.
3.2 RAG Pipeline
We implement a standard two-stage RAG framework with retrieval and generation phases. For retrieval, we use Contriever Izacard et al. (2022), a BERT-based dense retriever trained with unsupervised contrastive learning. Our knowledge corpus is Wikipedia-2018, preprocessed into 100-word non-overlapping passages. We encode these passages using Contriever and index them with FAISS [^1] https://github.com/facebookresearch/faiss for efficient retrieval. During training and development, the top-20 relevant documents are retrieved for each query. For question-answering, the top-5 retrieved documents are concatenated with the query as input for answer generation. The generator prompt is detailed in Appendix C.3 Table 3.
3.3 Robust RAG Training Setups
For our experiments, we evaluate two popular robust training settings:
3.3.1 Document Selection Strategies
We explore several document selection training strategies across three categories.
For basic QA capability evaluation, we use the Base Model in a few-shot setting, where the model performs inference without RAG-specific training.
For scenarios with relevant documents, we explore four strategies:
  • RALM: Fine-tunes the model by prepending golden retrieval text. Queries without golden documents in the top-20 are excluded, leading to reduced training set size.
  • RetRobust: Following Yoran et al. (2024), this method enhances robustness by randomly selecting top-ranked, low-ranked, or random passages during training.
  • Top-1 Document: Uses the highest-scoring retrieved document, which may not contain the correct answer, reflecting real-world retrieval challenges.
  • Golden Document: Selects the most relevant document containing the correct answer. If none exist in the top-20, it defaults to the top-1 document for consistency.
Model RAG Scenario HotpotQA NQ WebQuestions TriviaQA AVERAGE
EM F1 EM F1 EM F1 EM F1 EM F1
Llama-2-7b-chat-hf Base Model 3.30 12.34 1.21 10.61 0.00 13.08 4.32 20.27 2.21 14.08
RALM 26.21 36.42 32.17 42.68 33.81 45.85 50.28 60.17 35.62 46.28
RetRobust 31.29 43.65 37.71 49.49 36.33 47.98 57.61 67.52 40.74 52.16
Top-1 Doc 31.76 43.95 40.20 51.89 41.73 52.76 52.93 65.41 41.66 53.50
Golden Doc 30.67 42.78 36.50 47.77 39.93 52.11 50.25 63.28 39.34 51.49
[…]
Irrelevant Doc 31.01 42.98 37.08 48.93 39.21 50.79 51.97 64.70 39.82 51.85
[…]
IRM 34.38 47.11 40.96 53.07 53.96 61.62 57.58 69.08 46.72 57.72
$\Delta$ (Worst $\rightarrow$ Best) 16.19% 12.89% 21.87% 13.19% 48.53% 28.43% 13.55% 17.66% 15.95% 13.20%
Llama-3-8B-Instruct Base Model 23.31 32.60 30.04 41.59 26.98 43.25 58.80 66.45 34.78 45.97
RALM 27.64 38.10 35.19 46.10 47.84 56.98 54.75 63.27 41.36 51.11
RetRobust 36.06 48.99 43.28 55.04 52.88 62.10 59.06 67.77 47.82 58.48
Top-1 Doc 36.72 49.30 44.38 56.20 54.68 62.26 60.80 68.31 49.15 59.02
Golden Doc 35.52 48.31 41.35 53.13 48.92 58.41 58.26 66.99 46.01 56.71
[…]
Irrelevant Doc 35.31 47.92 42.45 54.41 46.76 57.67 58.97 66.57 45.87 56.64
[…]
IRM 35.19 48.08 41.13 53.14 53.96 61.64 57.15 69.13 46.86 58.00
$\Delta$ (Worst $\rightarrow$ Best) 14.04% 12.53% 7.90% 5.76% 16.94% 8.46% 11.74% 10.70% 10.92% 8.42%
Table 1: Performance comparison of different LLMs (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) across different robust RAG scenarios on four datasets (HotpotQA, NQ, WebQuestions, and TriviaQA). The row $\Delta$ (Worst $\rightarrow$ Best) indicates the performance gain achieved by the best method compared to the worst strategy among these training strategies, representing the benefit of sophisticated robust training methods.
To assess robustness against irrelevant information, we introduce two adversarial scenarios:
  • Random Document: Randomly selects a document from retrieved results, simulating unpredictable retrieval quality.
  • Irrelevant Document: Chooses a passage from another query’s retrieval results, ensuring no relevance to the current query.
3.3.2 Adversarial Loss Design
We assess two popular adversarial loss strategies:
  • RAAT. The regularization term in RAAT reduces the performance gap between the best and worst retrieval cases. By penalizing excessive performance disparity, RAAT ensures that the model remains stable even under challenging retrieval conditions, leading to improved robustness and generalization.
  • IRM. The regularization in IRM minimizes the variance in performance across different retrieval environments. By enforcing consistency, IRM mitigates sensitivity to distribution shifts, ensuring that the model performs reliably across diverse retrieval scenarios.
4 Does Sophisticated Robust Training Still Matter in Powerful Models?
To investigate whether sophisticated document selection strategies and adversarial loss designs are still essential for robust RAG performance as LLMs continue to evolve, we conduct comprehensive experiments across multiple LMs and datasets.
4.1 Do Sophisticated Document Selection Strategies Matter?
We conduct experiments to analyze the effectiveness of complex document selection strategies under Llama model families (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) in Table 1, and Qwen model families in Appendix Table 4.
Training with sophisticated documents enhances LM robustness for weak models
The experimental results presented in Tables 1 and 4 provide compelling evidence that robust training significantly improves model resilience when processing noisy documents. While base models exhibit substantial performance degradation when encountering noisy documents during inference, models that undergo robust training maintain consistent and superior QA performance across various document selection strategies. A notable example is the Llama-2-7b-chat-hf model’s performance (Table 1) on the HotpotQA dataset, where training with golden documents improves the EM score from 3.3 (Base Model) to 30.67 (Golden Doc), indicating increased resilience to document noise. This pattern of improvement is consistently observed across both Llama (Table 1) and Qwen (Table 4) model families, strongly indicating that robust training effectively mitigates the base models’ inherent vulnerability to noisy documents.
Training with random documents shows surprising effectiveness
We also notice that training with randomly selected documents exhibits remarkable effectiveness across all experimental configurations. Quantitative analysis shows that with Llama-2-7b-chat-hf, this approach achieves superior performance on WebQuestions compared to more sophisticated strategies. Similar observations emerge from experiments with Qwen1.5-7B-Chat, where random document selection achieves 46.04 EM on WebQuestions, approaching the optimal performance of 47.12 EM achieved by RetRobust. The consistency of these results across distinct model architectures suggests that the efficacy of random document selection represents an inherent characteristic of contemporary RAG systems.
Diminishing returns of sophisticated document selections as models evolve
Experimental results indicate that the performance gains from sophisticated document selection strategies diminish as models evolve. For instance, comparing Llama-2-7b-chat-hf with Llama-3-8B-Instruct, the improvement in performance due to advanced document selection strategies decreases significantly, with the $\Delta$ (Worst $\rightarrow$ Best) metric for NQ dropping from 21.87% to 7.90% EM. A similar trend is observed when comparing Qwen1.5-7B-Chat to Qwen2.5-7B-Instruct, where the performance improvement from sophisticated document selection also shows a noticeable reduction. These results suggest that as models become more advanced, their ability to process and utilize information improves independently of complex document selection strategies, leading to diminished returns from such methods.
4.2 Do Adversarial Loss Functions Matter?
To investigate whether the design of complex adversarial loss functions contributes to model performance, in Table 1 and 4, we also analyze the robustness of various adversarial loss designs.
Adversarial loss significantly enhances performance for weaker models
For the weaker model (Llama-2-7b-chat-hf), incorporating adversarial loss functions such as RAAT and IRM leads to a substantial improvement in performance compared to the base model or alternative document selection strategies. Specifically, while the base model achieves an average EM / F1 of only 2.21 / 14.08, applying adversarial loss functions boosts the performance to 40.64 / 50.99 for RAAT and 46.72 / 57.72 for IRM. This highlights the effectiveness of adversarial loss in improving model robustness to noisy documents, significantly enhancing both robustness and downstream inference performance. Notably, in some cases, RAAT and IRM outperform traditional document selection strategies (e.g., top-1 Doc and golden doc), demonstrating their value in scenarios where the model needs stronger guidance to handle noisy retrievals.
Figure 2: Comparison of F1 Scores on the HotpotQA dataset for training with Random Doc and Golden Doc across models with varying parameter sizes from 0.5B to 70B. We provide detailed results in Appendix Table 6.
Adversarial loss exhibits diminishing returns for stronger models
For the stronger models, such as Llama-3-8B-Instruct and Qwen2.5-7B-Instruct, the benefits of adversarial loss functions are less pronounced. The basic Llama-3-8B-Instruct already achieves an average EM / F1 of 34.78 / 45.97, and the introduction of adversarial losses (RAAT and IRM) results in only modest improvements, with average EM / F1 scores of 44.31 / 54.44 and 46.86 / 58.00, respectively. Similarly, Qwen2.5-7B-Instruct shows similar trends, with marginal gains from adversarial loss functions. These improvements are comparable to or even slightly worse than the performance achieved by random document selection (48.30 / 58.92) or top-1 document strategies (49.15 / 59.02). This indicates that the models’ inherent robustness reduces the impact of adversarial losses, and in some cases, may even hinder performance. We hypothesize that when the model’s internal robustness is already well-developed, additional constraints from adversarial losses may interfere with its ability to optimize on clean and relevant inputs.
Based on these findings, we conclude that both adversarial loss functions and document selection strategies are more beneficial for weaker models. For smaller models, these techniques significantly improve robustness by mitigating the impact of noisy documents. For stronger models with inherently robust performance, the advantages of complex loss designs and sophisticated document selection diminish, suggesting that simpler strategies like random document selection may be sufficient. This underscores the importance of tailoring training strategies to the model’s inherent capabilities.
4.3 Do Training Strategies Matter Across Model Scales?
Through analyzing Tables 1 and 4, we observe a counter-intuitive phenomenon: models trained with random documents outperformed those with golden documents, despite the latter containing ground truth answers typically yielding optimal results in standard SFT. To investigate whether this phenomenon extends beyond 7B-8B models, we conduct experiments across model scales from 0.5B to 70B parameters.
The results (detailed in Appendix Table 6) in Figure 2 demonstrate that for smaller models ($\leq$3B parameters), training with golden documents leads to superior performance. This suggests that smaller models, limited by their inherent capabilities, benefit more from high-quality golden documents containing direct answers. However, as model size increases, we observe that training with random documents becomes more effective. This shift can be attributed to larger models’ enhanced question-answering abilities and improved robustness. These models can better generalize to downstream tasks even when trained on random documents, which may contain noisier or less structured information. This finding indicates that sophisticated document selection strategies become less crucial as model size increases, revealing an important scaling property in model training.
5 Why Sophisticated Training No Longer Matters in Powerful Models?
In this section, we conduct comprehensive experiments to delve into the reasons why sophisticated robust training strategies may no longer be crucial in powerful models.
Figure 3: Confidence scores for correct and wrong answers on HotpotQA dataset, comparing Llama2 and Llama3 models across various robust training methods.
5.1 Powerful Models Enable Natural Calibration
To understand whether powerful models inherently possess the ability to distinguish reliable from unreliable answers, we take HotpotQA dataset as an example to examine the confidence calibration capabilities for Llama2 (Llama-2-7b-chat-hf) and Llama3 (Llama-3-8B-Instruct) models. Here, confidence is the mean of token-wise probabilities in the model’s generated answer, providing a measure of the model’s certainty in its predictions Varshney et al. (2023); Xiong et al. (2024). Figure 3 reveals striking differences in their calibration patterns. In the base model, Llama2 shows poor natural calibration levels, where confidence scores for incorrect answers (95.8) abnormally exceed those for correct ones (93.8). In contrast, Llama3 demonstrates inherently better calibration, maintaining higher confidence for correct answers (97.5) than incorrect ones (91.3) without any specialized complex training.
While robust training methods (Golden Doc, Random Doc, and IRM) can effectively calibrate confidence scores and improve the gap between correct and incorrect answers for Llama2 from -2 to 12 with IRM, the marginal benefits of these complex training strategies diminish as model architectures advance. For Llama3, which already achieves a 6.2 confidence gap naturally, the improvements from these training methods become less significant. This finding strongly suggests that advances in model architecture can effectively eliminate the need for complex robustness training procedures, as newer models come with better built-in calibration capabilities.
Figure 4: Generalization performance comparison across different strategies trained on HotpotQA (diagonal hatches bars) and evaluated on NQ, WebQuestions, TriviaQA datasets (plain bars). More results available in Appendix Table 8.
5.2 Simple Training Strategies Generalize Well in Powerful Models
We further investigate whether powerful models can maintain robust generalization across different datasets with simple training strategies. We fine-tune models on HotpotQA using four document selection approaches and evaluate their transfer performance on NQ, WebQuestions, and TriviaQA.
As shown in Figure 4, simple strategies demonstrate surprisingly strong generalization ability. Random document selection matches or even outperforms sophisticated IRM across all evaluation datasets, with performance gaps of less than 1%. For instance, in TriviaQA, random selection (69.5 F1) slightly surpasses both golden (68.2 F1) and IRM (68.7 F1) approaches. This trend becomes more pronounced in the powerful Llama-3-8B-Instruct, where the performance gap between simple and sophisticated strategies further narrows. The consistent cross-dataset performance, regardless of training strategy, indicates that model capacity, rather than training sophistication, is the key driver of generalization ability. These findings provide strong evidence that as models become more powerful, sophisticated training strategies become increasingly unnecessary.
Figure 5: Attention visualization for a QA case. Each subplot shows attention distribution heatmaps across different models, where cell ($i$, $j$) represents the average attention weight from the $j$-th attention layer to document $i$. Text highlighted in green indicates the correct answer and corresponding Doc1, blue indicates key terms from the query, and red indicates incorrect model predictions. The color intensity in the heatmaps indicates attention strength.
5.3 Powerful Models Learn Effective Attention Patterns with Simple Training
To provide a direct understanding of why simple training can achieve good performance, we visualize attention distributions across different training strategies. Figure 5 reveals that both sophisticated robust training methods (IRM) and simple approaches (random doc, top-1) achieve similar attention patterns, with clear focus on Doc1 (containing the correct answer) in middle layers (9-16). In contrast, the base model fails to attend to the correct document, generating a wrong answer. This finding provides direct evidence that powerful models can learn optimal attention mechanisms even with simple training strategies, making sophisticated training methods unnecessary.
Figure 6: Performance comparison with different numbers of random documents during training. Increasing the number of random documents consistently improves model performance.
Figure 7: Training curves comparison between random and golden document strategies using Llama-2-7b-chat-hf and LLama-3-8B-Instruct.
5.4 Training with Random Docs: Better Performance and Faster Convergence
We investigate why random training proves good performance from two aspects.
More random docs lead to better performance
We first vary the numbers of random documents (0 to 3) in training instances to examine how increasing random documents affects model performance. As shown in Figure 6, increasing random documents consistently improves F1 scores across both datasets. For Llama-2-7b-chat-hf, using 3 random documents (versus zero) improves F1 scores by 3 points on HotpotQA and 4 points on NQ. Llama-3-8B-Instruct shows similar gains, suggesting that powerful models can effectively learn from random documents, making sophisticated document selection relatively unnecessary.
Faster convergence with random training
The training dynamics in Figure 7 provide another evidence for why sophisticated document selection becomes unnecessary. Random document training not only achieves higher F1 scores (2-3 points improvement) but also reaches peak performance in fewer steps compared to golden document training. This faster convergence with better performance holds true for both model scales, indicating that simpler random training actually enables more efficient learning in powerful language models.
6 Insights and Future Directions
Our findings reveal a fundamental shift in RAG system: as models become more powerful, the marginal benefits of sophisticated training strategies diminish significantly. This observation has several important insights:
Simplified RAG architecture design
For powerful models, simple retrieval strategies (even random selection) can achieve comparable performance to sophisticated approaches. This enables streamlined RAG architectures by replacing elaborate document filtering mechanisms with simpler retrieval methods, substantially reducing system complexity without sacrificing performance.
Scalable RAG for open-domain tasks.
Larger models demonstrate robustness against noisy retrieval, suggesting that open-domain RAG systems can function effectively with minimal retrieval supervision. Instead of enforcing strict filtering of retrieved documents, future large-scale RAG systems can leverage weakly supervised learning, incorporating large-scale web data and noisy retrieval results to improve generalization.
Theoretical implications for model scaling.
This research reveals a previously unexplored aspect of scaling laws: the diminishing returns of complex training strategies as models grow larger. This challenges current theoretical frameworks and calls for new ones that better explain how training requirements evolve with model scale.
Broader impact on machine learning.
Our findings suggest that as models become more powerful, practitioners should prioritize architectural improvements and data quality over complex training strategies. This insight could lead to more efficient resource allocation in model development across various applications, from computer vision to natural language processing.
7 Conclusions
In this study, we systematically investigate whether complex robust training strategies remain necessary for RAG systems as model capacity grows. Our extensive experiments consistently show that while sophisticated training methods significantly enhance weaker models’ performance, their benefits diminish dramatically in more powerful models. Through the systematic analysis, we find that advanced models inherently possess strong confidence calibration, cross-dataset generalization, and effective attention patterns even with simple training. These findings suggest that RAG systems can benefit from simpler training strategies as models become more powerful, enabling more scalable applications with minimal complexity.
Limitations
While our study provides valuable insights into the impact of adversarial loss functions and document selection strategies on RAG model robustness, several limitations remain. First, our analysis is restricted to dense transformer-based models, leaving the effectiveness of these techniques on sparse models, such as mixture-of-experts (MoE) architectures, unexplored. Future work could investigate whether similar trends hold for sparsely activated models with dynamic routing mechanisms. Second, although we analyze the effectiveness of adversarial training, we do not explicitly examine its long-term stability or convergence properties, which may vary depending on hyperparameter choices and optimization dynamics. Additionally, while we demonstrate that stronger models exhibit diminishing returns from adversarial losses and document selection strategies, the precise mechanisms behind this phenomenon remain unclear. Further research is needed to understand how model capacity interacts with retrieval robustness.
Ethics Statements
Our study focuses on improving the robustness of RAG models, but several ethical considerations must be acknowledged. First, while adversarial training enhances model reliability, it does not eliminate the risk of biased or misleading outputs, particularly when retrieval sources contain inherent biases or misinformation. Future work should explore fairness-aware adversarial training to mitigate potential harms. Second, our findings suggest that stronger models require less intervention in document selection and loss design, which may influence resource allocation in real-world applications. Researchers and practitioners should ensure that model improvements do not disproportionately benefit well-resourced institutions while leaving smaller models less robust. Lastly, our experiments are conducted on widely used benchmark datasets, which may not fully reflect the diversity of real-world information needs. We encourage further research on robustness evaluation across varied domains, including low-resource languages and specialized knowledge fields, to ensure equitable advancements in RAG technology.
Appendix A Related Work
A.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) has been developed to enhance the reasoning and generation capabilities of LLMs by integrating external information. RAG improves the overall performance of LLMs by retrieving relevant documents that provide additional context and support for the generation process. However, dense retrievers are not always perfect and can sometimes recall irrelevant or erroneous documents, which can degrade performance Jiang et al. (2023); Deng et al. (2023); Ding et al. (2024). To mitigate this issue, adaptive retrieval techniques and robust RAG training methods have been proposed. Adaptive retrieval techniques retrieve external information only when LLMs encounter unknowns, reducing the risk of integrating misleading or incorrect information Jiang et al. (2023); Mallen et al. (2023); Asai et al. (2023). Additionally, robust training focuses on enhancing LLMs’ response to various retrieval noises, aiming to efficiently obtain adversarial examples that improve model robustness while reducing training overhead Fang et al. (2024); Zhu et al. (2024a); Xiang et al. (2024).
A.2 Robust Training for RAG Systems
The performance of LLMs in RAG systems can be compromised by irrelevant context Cuconasu et al. (2024); Wu et al. (2024). Recently, researchers have focused on this issue and proposed effective adversarial training methods to enhance the robustness of LLMs against noisy documents. RetRobust Yoran et al. (2024) suggests using a mix of relevant and irrelevant documents to train LLMs to withstand noisy contexts. RAAT Fang et al. (2024) and ATM Zhu et al. (2024a) employ adversarial training to dynamically adjust the model’s training process, aiming to maintain robustness against noise and generate accurate answers. RobustRAG Xiang et al. (2024) proposes to enhances RAG model robustness against retrieval corruption attacks through an isolate-then-aggregate strategy to achieve certifiable robustness. Additionally, there is a body of work focused on identifying key documents from noisy contexts to facilitate effective question answering through knowledge rewriting, refinement, or filtering Zhu et al. (2024b); Qiao et al. (2024); Jin et al. (2024); Zeng et al. (2024). However, these robust training approaches are primarily applied to small or weak LMs with fewer than 7 billion parameters. Thus, there’s an urgent need to explore whether complex robust training is still necessary to improve the robustness and generalization of bigger or stronger models when dealing with noisy contexts.
Appendix B Robustness Training for RAG
B.1 Retrieval-augmented Adaptive Adversarial Training
To improve the robustness of retrieval-augmented language models (RALMs) against retrieval noise, RAAT Fang et al. (2024) incorporates an adversarial loss into the standard supervised fine-tuning (SFT). The model processes one golden context and three adversarial samples at each iteration, optimizing for the most challenging perturbation. The adversarial objective follows a min-max strategy:
$\mathcal{L}_{\text{max}}=\max_{da\in DA}\mathcal{L}(\theta,da(q),a),$ (2)
where $\mathcal{L}$ represents the generation loss function to noisy retrievals, and $q^{\prime}=da(q)$ represents the augmented noise context of $q$. A regularization term controls excessive sensitivity to noise:
$\mathcal{L}_{\text{ada}}=\mathcal{L}_{\text{max}}+w_{\text{reg}}\cdot|\mathcal{L}_{\text{max}}-\mathcal{L}_{\text{min}}|_{2}^{2},$ (3)
The adversarial loss $\mathcal{L}_{\text{ada}}$ combines both terms, where the regularization term helps stabilize training by mitigating excessive sensitivity to retrieval noise. The regularization term, calculated as the square of the difference between $\mathcal{L}_{\text{max}}$ and $\mathcal{L}_{\text{min}}$, encourages a more balanced optimization, preventing the model from overreacting to the most challenging perturbations.
B.2 Multiagent Iterative Tuning Optimization
ATM Zhu et al. (2024a) steers the generator to have a robust perspective of useful documents for question answering with the help of an auxiliary attacker agent. The generator is trained to maximize answer correctness while minimizing its sensitivity to adversarial perturbations. It receives the user query along with a document list, which may contain both relevant and fabricated information introduced by the attacker. Its objective is to generate accurate responses as long as sufficient truthful information is present while reducing the impact of misleading or fabricated content. To achieve this, the generator learns to identify and leverage relevant documents while ignoring noisy ones, regardless of whether they originate from the original retrieval $\mathcal{D}$ or adversarial perturbations $\mathcal{D}^{\prime}$. This can be formalized as maximizing the objective:
$G(a\mid q,\mathcal{D}^{\prime})-\mathrm{dist}\left[G(a\mid q,\mathcal{D}),G(a\mid q,\mathcal{D}^{\prime})\right],$ (4)
where ${G}(\cdot)$ represents the language model probability of generating an answer, and $\mathrm{dist}\left[\cdot\right]$ measures the divergence between outputs under different document conditions. The generator and the attacker are tuned adversarially for several iterations. After rounds of multi-agent iterative tuning, the generator can eventually better discriminate useful documents amongst fabrications.
B.3 Invariant Risk Minimization
Invariant Risk Minimization (IRM) aims to learn representations that remain stable across different environments, improving generalization under distribution shifts. To enhance the robustness of retrieval-augmented generation (RAG) across varying retrieval conditions, the V-REx objective Krueger et al. (2021) can also be adapted to enforce risk invariance across different retrieval environments. Given retrieval environments =$\mathcal{E}=\{1,\dots,m\}$, where each $e\in\mathcal{E}$ corresponds to a specific retrieval scenario (e.g., golden documents, top-$k$ retrieved, noisy retrieval), we define the empirical risk $\mathcal{R}_{e}(\theta)$ of the model parameterized by $\theta$. The training objective is formulated as:
$\mathcal{R}_{\textrm{V-REx-RAG}}(\theta)=$ $\beta\;\mathrm{Var}(\{\mathcal{R}_{1}(\theta),...,\mathcal{R}_{m}(\theta)\})$ (5)
$+\sum^{m}_{e=1}\mathcal{R}_{e}(\theta),$
where $\beta\geq 0$ controls the trade-off between minimizing average risk and enforcing risk invariance across retrieval environments. A higher $\beta$ reduces performance discrepancies caused by retrieval variations, improving generalization under both high-quality and noisy retrieval conditions.
Appendix C Appendix
C.1 Dataset Statistics
Detailed statistics of these datasets we used are listed in Table 2.
Dataset Type # Train # Dev
NQ single-hop 79,168 8,757
WebQuestions single-hop 2,474 278
TriviaQA multi-hop 78,785 8,837
HotpotQA multi-hop 90,447 7,405
Table 2: Statistics of different datasets
C.2 Implementation Details
For model training, we utilize the LLaMA-Factory library to facilitate efficient LLM finetuning. We train our models with a learning rate of 1e-6 and set the maximum number of epochs to 3. To optimize the training process, we employ bfloat16 precision and DeepSpeed ZeRO-3 for distributed training. The gradient accumulation steps are set to 2 with a batch size of 8. For inference, we leverage vLLM to ensure efficient model serving. We maintain the default decoding parameters for each model during inference, including the default temperature and top-p sampling probabilities, to ensure fair comparison across different model variants. All training and inference procedures are conducted on 8x NVIDIA A100 80GB GPUs.
C.3 Prompt Instruction
We present the prompt used to guide LM inference on noisy contexts in Table 3. Specifically, we concatenate the top-5 retrieved documents into a single passage, append the input question afterward, and prompt the LM to generate a concise answer.
You need to answer my question and complete the question-and-answer pair following the format provided in the example. The answers should be short phrases or entities, not full sentences. Here are some examples to guide you.
Example 1:
Question: What is the capital of France?
Answer: Paris
Example 2:
Question: Who invented the telephone?
Answer: Alexander Graham Bell
Example 3:
Question: Which element has the atomic number 1?
Answer: Hydrogen
### Retrieved Documents: {}
### Question: {}
### Answer:
Table 3: The generator prompt we used in our experiments.
Model RAG Scenario HotpotQA NQ WebQuestions TriviaQA AVERAGE
EM F1 EM F1 EM F1 EM F1 EM F1
Qwen1.5-7B-Chat Base Model 24.15 34.06 24.31 35.27 23.02 38.65 48.43 57.65 29.98 41.41
RALM 24.44 34.23 29.83 40.22 44.24 53.46 46.40 55.96 36.23 45.97
RetRobust 29.05 40.99 34.82 45.92 47.12 55.39 49.16 59.01 40.04 50.33
Top-1 Doc 29.68 41.57 34.16 45.44 43.88 52.71 49.61 59.03 39.33 49.69
Golden Doc 28.66 40.55 32.93 43.99 43.88 52.74 49.25 58.87 38.68 49.04
[…]
Irrelevant Doc 28.12 39.87 32.74 43.94 43.88 52.67 48.33 58.18 38.27 48.67
[…]
$\Delta$ (Worst $\rightarrow$ Best) 21.44% 21.44% 16.73% 14.17% 9.02% 5.71% 8.99% 5.49% 10.52% 9.48%
Qwen2.5-7B-Instruct Base Model 25.10 35.02 26.97 38.15 25.90 42.46 53.86 62.73 32.96 44.59
RALM 26.37 36.30 31.43 42.00 41.73 51.98 53.92 62.74 38.36 48.26
RetRobust 30.47 42.26 34.60 46.00 45.68 53.80 54.43 63.63 41.30 51.42
Top-1 Doc 30.24 42.61 34.72 46.20 46.04 54.13 53.65 63.01 41.16 51.49
Golden Doc 30.28 41.84 33.77 45.11 44.60 53.18 54.35 63.55 40.75 50.92
[…]
Irrelevant Doc 29.76 41.67 33.64 44.87 42.09 52.72 53.38 63.01 39.72 50.57
[…]
$\Delta$ (Worst $\rightarrow$ Best) 15.55% 17.38% 10.47% 10.00% 11.19% 5.43% 1.97% 2.42% 7.64% 6.70%
Table 4: Performance comparison of different LLMs (Qwen1.5-7B-Chat and Qwen2.5-7B-Instruct) across different robust RAG scenarios on four datasets (HotpotQA, NQ, WebQuestions, and TriviaQA). The row $\Delta$ (Worst $\rightarrow$ Best) represents the performance gain achieved by the best method compared to the worst strategy among these training strategies, highlighting the benefit of sophisticated robust training methods.
Model Document Distribution HotpotQA NQ WebQuestion TriviaQA AVERAGE
EM F1 EM F1 EM F1 EM F1 EM F1
Llama-2-7b-chat-hf Base Model 3.30 12.34 1.21 10.61 0.00 13.08 4.32 20.27 2.21 14.08
1 Golden Doc 30.67 42.78 36.50 47.77 39.93 52.11 50.25 63.28 39.34 51.49
1 Random Doc 30.94 43.11 38.16 49.78 42.45 53.97 52.72 65.52 41.07 53.10
0 Random + 3 Golden 31.79 44.03 36.68 47.90 43.02 53.43 56.55 64.97 42.01 52.58
1 Random + 2 Golden 32.82 44.77 36.70 48.21 43.38 54.34 56.93 65.67 42.46 53.25
2 Random + 1 Golden 33.32 45.45 37.07 48.67 44.10 54.82 57.56 65.90 43.01 53.71
3 Random + 0 Golden 34.74 46.79 37.35 49.34 45.18 55.70 57.77 66.40 43.76 54.56
Llama-3-8B-Instruct Base Model 23.31 32.60 30.04 41.59 26.98 43.25 58.80 66.45 34.78 45.97
1 Golden Doc 35.52 48.31 41.35 53.13 48.92 58.41 58.26 66.99 46.01 56.71
1 Random Doc 35.98 49.05 43.37 55.43 53.24 62.55 60.62 68.64 48.30 58.92
0 Random + 3 Golden 35.96 49.01 40.88 52.91 50.25 59.13 54.95 66.48 45.51 56.88
1 Random + 2 Golden 36.22 49.07 41.22 53.13 50.41 59.97 55.27 66.77 45.78 57.24
2 Random + 1 Golden 36.77 49.59 41.75 53.48 51.49 61.02 55.39 66.74 46.35 57.71
3 Random + 0 Golden 37.33 50.30 42.45 54.32 52.05 61.16 56.20 67.83 47.01 58.40
Table 5: Experimental comparison of RAG robust training using different numbers of random documents and golden documents mixed together, with Llama-2-7b-chat-hf and Llama-3-8B-Instruct models across four datasets.
Model RAG Scenario HotpotQA NQ
EM F1 EM F1
Qwen2.5-0.5B-Instruct RetRobust 17.69 26.63 17.16 25.44
Top-1 Doc 18.28 27.38 18.00 26.52
Golden Doc 19.11 28.04 19.32 27.49
Random Doc 18.10 27.23 16.27 24.62
Qwen2.5-1.5B-Instruct RetRobust 23.35 34.13 25.01 35.35
Top-1 Doc 24.08 34.31 25.53 35.75
Golden Doc 23.70 34.03 25.61 35.52
Random Doc 23.50 33.68 24.63 34.89
Qwen2.5-3B-Instruct RetRobust 26.29 37.34 29.46 40.38
Top-1 Doc 26.17 37.33 29.86 40.67
Golden Doc 26.17 37.16 30.08 40.41
Random Doc 26.28 37.26 29.34 40.09
Mistral-7B-Instruct-v0.2 RetRobust 35.92 49.03 43.19 55.51
Top-1 Doc 35.80 49.08 44.35 56.29
Golden Doc 34.67 47.46 41.52 53.41
Random Doc 35.52 48.71 43.87 55.84
Qwen2.5-32B-Instruct RetRobust 33.96 46.28 35.67 48.77
Top-1 Doc 34.47 46.88 35.90 48.80
Golden Doc 33.68 45.62 35.29 47.69
Random Doc 33.81 46.31 36.09 49.03
Llama-2-70b-chat-hf RetRobust 36.22 49.58 41.49 54.73
Top-1 Doc 35.45 48.75 40.39 53.92
Golden Doc 34.27 47.72 39.23 51.69
Random Doc 34.90 48.31 40.39 53.92
Llama-3-70B-Instruct RetRobust 40.78 54.67 45.84 59.25
Top-1 Doc 40.61 55.37 46.05 59.61
Golden Doc 40.65 54.15 44.80 57.85
Random Doc 39.91 54.71 43.71 57.86
Qwen2.5-72B-Instruct RetRobust 37.06 49.96 39.99 52.70
Top-1 Doc 37.52 50.30 39.87 52.81
Golden Doc 36.73 49.66 38.86 51.73
Random Doc 36.75 50.05 39.95 52.86
Table 6: Performance comparison of LLM with different parameter sizes (from 0.5B to 72B) in different robust RAG scenarios on HotpotQA and NQ datasets.
Models RAG Scenario HotpotQA NQ WebQ TriviaQA
Correct Wrong Correct Wrong Correct Wrong Correct Wrong
Llama2-7b-chat-hf Base Model 93.80 95.82 95.25 95.87 0.00 96.21 94.56 96.04
Golden Doc 96.33 86.80 96.28 87.78 97.11 89.42 89.69 79.29
Random Doc 96.45 84.54 96.06 85.58 96.13 89.00 89.22 78.60
IRM 95.78 83.33 95.77 85.09 97.06 91.67 92.96 80.47
Llama-3-8B-Instruct Base Model 97.53 91.30 97.59 91.30 97.90 89.22 98.15 90.08
Golden Doc 96.78 85.82 96.36 86.48 94.86 85.89 93.03 76.31
Random Doc 96.46 83.63 96.15 84.50 94.71 84.83 96.08 81.97
IRM 95.84 80.68 95.55 82.33 98.63 93.16 89.31 76.55
Qwen1.5-7B-Chat Base Model 98.00 93.66 98.50 94.03 97.64 92.34 98.00 92.80
Golden Doc 96.33 83.87 95.94 84.39 94.20 84.82 92.90 76.17
Random Doc 95.55 80.39 95.13 80.39 93.68 82.59 95.15 83.18
IRM 95.07 78.05 93.99 77.91 98.41 91.10 89.87 72.35
Qwen2.5-7B-Instruct Base Model 98.52 94.54 98.98 95.42 97.15 95.54 98.62 93.80
Golden Doc 96.25 83.91 96.00 84.83 95.33 87.23 96.24 82.33
Random Doc 95.92 80.28 95.57 80.46 95.28 84.24 96.51 82.87
IRM 96.04 80.20 95.31 80.68 97.75 91.73 93.78 78.70
Table 7: Comparison of answer confidence scores for different models after robust training strategies on HotpotQA, NQ, WebQ and TriviaQA datasets.
Model RAG Scenario HotpotQA NQ WebQuestions TriviaQA
EM F1 EM F1 EM F1 EM F1
Llama-2-7b-chat-hf Base Model 3.30 12.34 1.21 10.61 0.00 13.08 4.32 20.27
RALM 26.21 36.42 27.90 38.41 30.22 46.22 55.36 63.44
RetRobust 31.29 43.65 30.72 43.69 30.94 48.52 59.56 68.58
Top-1 Doc 31.76 43.95 31.31 44.02 30.94 49.73 59.99 68.67
Golden Doc 30.67 42.78 30.59 42.52 30.22 46.58 58.59 67.25
Random Doc 30.94 43.11 31.08 43.92 30.58 48.58 60.42 69.16
Irrelevant Doc 31.01 42.98 31.10 43.41 29.14 46.70 59.25 67.92
IRM 34.38 47.11 30.79 43.06 31.29 49.51 59.57 68.26
Llama-3-8B-Instruct Base Model 23.31 32.60 30.04 41.59 26.98 43.25 58.80 66.45
RALM 27.64 38.10 30.34 41.25 30.22 44.94 58.72 66.90
RetRobust 36.06 48.99 34.05 47.10 31.65 49.53 61.62 70.99
Top-1 Doc 36.72 49.30 34.32 47.12 30.22 49.96 62.16 71.28
Golden Doc 35.52 48.31 33.78 46.45 29.86 47.26 60.72 69.89
Random Doc 35.98 49.05 34.13 47.30 28.70 47.94 61.94 71.16
Irrelevant Doc 35.31 47.92 33.80 47.03 31.65 50.63 61.37 70.77
IRM 35.19 48.08 34.29 47.15 30.94 49.09 61.68 70.80
Qwen1.5-7B-Chat Base Model 24.15 34.06 24.31 35.27 23.02 38.65 48.43 57.65
RALM 24.44 34.23 25.00 35.02 23.74 37.56 49.97 58.42
RetRobust 29.05 40.99 25.84 37.67 27.34 43.27 52.97 61.93
Top-1 Doc 29.68 41.57 25.21 37.12 25.18 42.39 52.27 61.90
Golden Doc 28.66 40.55 25.43 36.84 24.46 41.78 51.26 60.34
Random Doc 28.83 40.66 25.88 38.04 26.62 43.43 52.47 61.78
Irrelevant Doc 28.12 39.87 24.67 36.46 25.54 42.16 51.45 60.55
IRM 26.95 38.67 25.98 37.63 25.54 41.77 52.02 61.37
Qwen2.5-7B-Instruct Base Model 25.10 35.02 26.97 38.15 25.90 42.46 53.86 62.73
RALM 26.37 36.30 26.98 37.68 26.98 42.99 54.43 62.86
RetRobust 30.47 42.26 28.58 41.25 28.78 45.61 58.52 67.21
Top-1 Doc 30.24 42.61 28.85 41.41 28.78 46.31 58.74 67.34
Golden Doc 30.28 41.84 28.93 41.10 28.78 46.68 57.64 66.39
Random Doc 30.25 42.22 28.73 41.19 28.42 44.72 58.40 67.38
Irrelevant Doc 29.76 41.67 28.40 41.00 27.70 43.90 58.40 67.38
IRM 29.58 41.73 28.77 41.47 25.54 42.84 58.28 67.17
Table 8: Comparison of different document selection strategies on three datasets (NQ, WebQuestions, TriviaQA) after training on HotpotQA. The gray cells indicate results evaluated on HotpotQA, while the orange cells indicate results evaluated on other datasets to assess generalization ability.

### v2

@@
-Revisiting Robust RAG: Do We Still Need Complex Robust Training in the Era of Powerful LLMs?
-Abstract
-Retrieval-augmented generation (RAG) systems often suffer from performance degradation when encountering noisy or irrelevant documents, driving researchers to develop sophisticated training strategies to enhance their robustness against such retrieval noise.
-However, as large language models (LLMs) continue to advance, the necessity of these complex training methods is increasingly questioned.
-In this paper, we systematically investigate whether complex robust training strategies remain necessary as model capacity grows.
-Through comprehensive experiments spanning multiple model architectures and parameter scales, we evaluate various document selection methods and adversarial training techniques across diverse datasets.
-Our extensive experiments consistently demonstrate that as models become more powerful, the performance gains brought by complex robust training methods drop off dramatically.
-We delve into the rationale and find that more powerful models inherently exhibit superior confidence calibration, better generalization across datasets (even when trained with randomly selected documents), and optimal attention mechanisms learned with simpler strategies.
-Our findings suggest that RAG systems can benefit from simpler architectures and training strategies as models become more powerful, enabling more scalable applications with minimal complexity.
-Revisiting Robust RAG: Do We Still Need Complex Robust Training
-in the Era of Powerful LLMs?
-Hanxing Ding1,2thanks:   Equal contributions.
-Shuchang Tao11footnotemark: 1  Liang Pang1thanks:   Corresponding author Zihao Wei1,2 Liwei Chen3Kun XuHuawei Shen1,2 Xueqi Cheng1,2 1Key Laboratory of AI Safety, Institute of Computing Technology, Chinese Academy of Sciences 2 University of Chinese Academy of Sciences  3 Kuaishou Technology {dinghanxing18s, pangliang, weizihao22z, shenhuawei, cxq}@ict.ac.cn
-"Entities should not be multiplied unnecessarily."
-— Occam’s razor
-1 Introduction
-Modern LLMs excel in various NLP tasks, such as text generation, knowledge reasoning, and question answering Ye et al.
-(2024); Tao et al.
-(2024); Zhao et al.
-(2023).
-Their contextual relevance and coherence make them valuable tools across domains.
-Despite their broad applications, LLMs struggle with external knowledge integration, particularly in RAG systems where poor retrieval results can lead to inaccurate or misleading responses Xu et al.
-(2024); Tan et al.
-(2024); Ding et al.
-(2024).
-This issue exposes a key limitation in the robustness of LLMs with imperfect retrieval results.
-Figure 1: Comparison of EM scores of different document selection strategies on the TriviaQA dataset.
-As the model capabilities increase, the performance gap ($\Delta$) between the best and worst strategies among the four training methods decreases from 14.65% to 4.36%.
-To address this challenge, researchers have proposed various robust training methods to enhance RAG’s resilience against noisy contexts.
-Approaches mainly focus on selecting high-quality documents or incorporating additional adversarial loss regularization terms during training Yoran et al.
-(2024); Fang et al.
-(2024); Krueger et al.
-(2021); Zhu et al.
-(2024a); Jin et al.
-(2024); Zeng et al.
-(2024).
-For instance, RetRobust Yoran et al.
-(2024) suggests using a strategic mixture of relevant and irrelevant documents to train LLMs to withstand noisy contexts.
-RAAT Fang et al.
-(2024) and ATM Zhu et al.
-(2024a) introduce adversarial regularization terms to encourage consistent model performance across different document contexts, making the model less sensitive to retrieval noise.
-Given the increasing capabilities of LLMs, a fundamental question emerges: do sophisticated robust training strategies remain necessary as model capacity grows?
-We first conduct a simple preliminary experiment to investigate this inquiry.
-As shown in Figure 1, we compare the EM scores of four different training strategies on the TriviaQA dataset.
-Notably, the performance gap ($\Delta$) between the best and worst strategies shrinks dramatically from 14.65% to 4.36% as the model capacity increases.
-This significant diminishing gap suggests that powerful LLMs might naturally possess enhanced robustness, potentially challenging the necessity of complex robust training techniques.
-In this paper, we present a systematic investigation into the necessity of sophisticated robust training strategies as model capabilities advance.
-Through extensive experiments across various language models with different capabilities levels, we uncover a counterintuitive phenomenon: while models with limited capabilities benefit from high-quality document selection and complex adversarial losses, these benefits diminish substantially as model capability increases.
-Notably, for highly capable models, training with randomly selected or even seemingly irrelevant documents can achieve comparable or superior performance to sophisticated training strategies.
-We further conduct an in-depth study on this counter-intuitive phenomenon through a systematic analysis.
-Through inherent confidence calibration analysis, we discover that newer powerful models naturally exhibit superior discrimination between correct and incorrect responses.
-Moving to external performance, we find these models demonstrate remarkable cross-dataset generalization with basic random document training.
-Our visualization of attention patterns further reveals the underlying mechanism: powerful models inherently learn effective attention allocation even with simple training.
-These findings, from internal capabilities to external performance to mechanism analysis, indicate that powerful models may naturally possess fundamental desired capabilities, potentially reducing the need for complex training strategies.
-We provide insights for future RAG development, suggesting: simplified architecture design principles for powerful models, opportunities for scalable open-domain applications with minimal supervision, and theoretical perspectives on how training requirements evolve with model scaling.
+On the Diminishing Returns of Complex Robust RAG Training in the Era of Powerful LLMs
+Abstract.
+Retrieval-augmented generation (RAG) systems traditionally employ sophisticated training strategies to enhance robustness against retrieval noise.
+In this work, we investigate a critical question: does the benefit of these complex robust training methods diminish as language models become more powerful?
+Through systematic evaluation across multiple model scales and question-answering datasets, our analysis reveals a consistent trend: the marginal robustness benefit of sophisticated training strategies decreases substantially as model capacity increases.
+While smaller models show significant performance improvements from complex document selection and adversarial objectives, more capable models achieve comparable or even superior performance with simpler training approaches.
+Further investigation demonstrates that stronger models naturally exhibit better confidence calibration, cross-dataset generalization capability, and more effective attention patterns, even under simple training regimes.
+These findings suggest that as foundation models evolve, the engineering effort invested in complex robust training may yield diminishing returns, indicating that simplified RAG pipelines could suffice for powerful models while maintaining competitive performance.
+Large Language Models, Retrieval-augmented Generation, Robust Training
+journalyear: 2025copyright: ccconference: Proceedings of the 2025 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region; December 7–10, 2025; Xi’an, Chinabooktitle: Proceedings of the 2025 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region (SIGIR-AP 2025), December 7–10, 2025, Xi’an, Chinadoi: 10.1145/3767695.3769518isbn: 979-8-4007-2218-9/2025/12ccs: Information systems Question answering
+1. Introduction
+Modern LLMs excel in various NLP tasks including text generation, knowledge reasoning, and question answering (Ye et al., 2024; Tao et al., 2024; Zhao et al., 2023), yet they struggle with external knowledge integration, particularly in RAG systems where poor retrieval can lead to inaccurate responses (Xu et al., 2024; Tan et al., 2024; Ding et al., 2024).
+To address this challenge, researchers have proposed robust training methods focusing on high-quality document selection or adversarial loss regularization (Yoran et al., 2024; Fang et al., 2024; Krueger et al., 2021; Zhu et al., 2024b; Jin et al., 2024; Zeng et al., 2024).
+For instance, RetRobust (Yoran et al., 2024) uses a mixture of relevant and irrelevant documents during training, while RAAT (Fang et al., 2024) and ATM (Zhu et al., 2024b) incorporate adversarial regularization terms to maintain consistent performance across varied document contexts.
+Figure 1.
+Performance comparison of different robust training strategies on the TriviaQA dataset.
+As model capabilities increase, the marginal robustness benefit ($\Delta$), i.e., performance gain between the best and worst strategies among the four training methods decreases from 14.65% to 4.36%.
+With the rapid advancement of LLM capabilities, a critical research question arises: how does the effectiveness of sophisticated robust training strategies change as model capacity grows?
+To investigate this question empirically, we conduct a preliminary analysis comparing four distinct training strategies on the TriviaQA dataset, as illustrated in Figure 1.
+Our results reveal a striking pattern: the performance gain $\Delta$ between the best and worst strategies decreases from 14.65% to 4.36% as model capacity scales up.
+This pronounced reduction in $\Delta$ suggests a law of diminishing returns, wherein the marginal benefit of complex training methods becomes increasingly negligible as model capacity expands.
+In this paper, building upon this observation, we conduct a systematic investigation to validate and characterize the diminishing returns of robust RAG training as model capacity increases.
+We propose an evaluation framework that quantifies the marginal robustness benefit ($\Delta$)—the performance gain introduced by sophisticated document selection and adversarial training methods.
+Our comprehensive experiments confirm and strengthen the initial finding: as model capability increases, the marginal robustness benefit of complex training diminishes sharply.
+While smaller models significantly benefit from curated documents and adversarial losses, stronger models achieve comparable or even superior performance with simpler setups.
+Strikingly, training high-capacity models with randomly selected documents often matches or exceeds the performance of advanced robust training methods, challenging fundamental assumptions about robust training strategies.
+To uncover the mechanisms behind this phenomenon, we perform a comprehensive analysis of model behavior.
+Our findings show that stronger models possess inherent capabilities that reduce reliance on complex training.
+They exhibit superior confidence calibration, accurately distinguishing correct from incorrect outputs without additional supervision.
+Even when trained with random documents, these models generalize well across datasets.
+Attention visualizations further reveal that high-capacity models naturally learn effective attention patterns under simple training, explaining the diminishing gains from sophisticated methods.
+These findings reveal a shift in how robust RAG systems should be designed: as model capacity increases, simpler training and retrieval strategies become not only sufficient but preferable.
+This shift opens up new opportunities for scalable RAG applications, and prompts a rethinking of training priorities in the context of model scaling.
 The main contributions of this work are:
-  • Finding: We reveal that as model capabilities increase, the benefits of sophisticated training strategies diminish substantially.
-  • Rationale: Through systematic analysis, we trace this finding to powerful models’ inherent capabilities, demonstrating that they naturally possess strong confidence calibration, generalization, and effective attention mechanisms, regardless of training sophistication.
-  • Insights: We provide practical guidelines for future RAG development, advocating for architectural simplification with powerful models and highlighting new theoretical perspectives on model scaling laws.
-2 Background
-2.1 Retrieval-Augmented Generation
-Retrieval-Augmented Generation (RAG) enhances language model performance by incorporating external knowledge retrieval into the generation process.
-Given a query $q$, a retriever $\mathcal{R}$ selects a set of $k$ relevant document sets $\mathcal{D}=\{d_{1},d_{2},\dots,d_{k}\}$ from a large corpus, providing additional factual grounding for the model.
-These retrieved documents serve as context to improve the model’s reasoning and factual accuracy.
-The language model $G$ then generates an answer $a$ based on the augmented input, following the conditional probability formulation:
-$G(a\mid q,\mathcal{D})=\prod_{t=1}^{T}P(a_{t}\mid a_{<t},q,\mathcal{D};\theta),$ (1)
-where $a=(a_{1},a_{2},\dots,a_{T})$ represents the generated sequence, and $\theta$ denotes the model parameters.
-The retrieval-augmented setup enables the model to dynamically condition its responses on external information, improving its ability to generate factually consistent answers.
-However, RAG models remain susceptible to retrieval noise, such as irrelevant or misleading documents, which can degrade performance and lead to hallucinations.
-2.2 Robust Training for RAG
-Document Selection Strategy
-A common strategy for robust training in RAG models is to select high-quality documents from either the golden set or top-ranked passages for fine-tuning.
-Traditionally, supervised fine-tuning (SFT) relies on high-quality golden data, where each query is paired with a carefully curated document containing the correct answer.
-This ensures that the model learns from reliable sources, leading to improved factual accuracy.
-Naturally, leveraging such golden documents in RAG training is a straightforward and effective approach.
-However, in real-world scenarios, the golden document may not always be retrieved due to limitations in the retriever.
-In such cases, training only on golden passages lead to unrealistic expectations during inference.
-A more practical approach is to incorporate top-1 retrieved documents, which better reflect what the model will encounter at test time.
-Even if these documents are imperfect, training on them enhances the model’s adaptability.
-Further, drawing from a mixture of top-ranked, lower-ranked, or randomly sampled passages, as suggested by RetRobust Yoran et al.
-(2024), exposes the model to diverse retrieval conditions, improving its robustness to noisy or suboptimal results.
-This strategy ensures that RAG models generalize better across different retrieval settings, making them more reliable in real-world applications.
-Adversarial Loss Design
-To enhance the robustness of RAG models against retrieval variations, robust training for RAG introduces a general training framework that combines standard SFT with a robustness-enforcing regularization term.
-During training, the model learns across multiple retrieval environments, which may include golden documents, top-k retrieved results, or adversarially perturbed documents.
-The overall training objective consists of two components: the first part is the standard training loss across different retrieval environments, and the second part is a regularization term designed to reduce model performance sensitivity to retrieval variations.
-Different forms of regularization strategies are detailed in Appendix B.
-3 Experimental Setups
-In this section, we introduce the dataset and evaluation metrics we used to assess the robustness performance of various robust training strategies.
-3.1 Datasets and Evaluation Metrics
-For our experiment, we evaluate on four widely-used question answering datasets: (1) single-hop QA, including NaturalQuestions (NQ) Kwiatkowski et al.
-(2019) and WebQuestions Berant et al.
-(2013); and (2) multi-hop QA, including TriviaQA Joshi et al.
-(2017) and HotpotQA Yang et al.
-(2018).
-All experimental results are evaluated on their development splits using the Exact Match (EM) and F1 metrics.
-Detailed statistics of these datasets are listed in Appendix C.1 Table 2.
-3.2 RAG Pipeline
-We implement a standard two-stage RAG framework with retrieval and generation phases.
-For retrieval, we use Contriever Izacard et al.
-(2022), a BERT-based dense retriever trained with unsupervised contrastive learning.
-Our knowledge corpus is Wikipedia-2018, preprocessed into 100-word non-overlapping passages.
-We encode these passages using Contriever and index them with FAISS [^1] https://github.com/facebookresearch/faiss for efficient retrieval.
-During training and development, the top-20 relevant documents are retrieved for each query.
-For question-answering, the top-5 retrieved documents are concatenated with the query as input for answer generation.
-The generator prompt is detailed in Appendix C.3 Table 3.
-3.3 Robust RAG Training Setups
-For our experiments, we evaluate two popular robust training settings:
-3.3.1 Document Selection Strategies
-We explore several document selection training strategies across three categories.
-For basic QA capability evaluation, we use the Base Model in a few-shot setting, where the model performs inference without RAG-specific training.
-For scenarios with relevant documents, we explore four strategies:
-  • RALM: Fine-tunes the model by prepending golden retrieval text. Queries without golden documents in the top-20 are excluded, leading to reduced training set size.
-  • RetRobust: Following Yoran et al. (2024), this method enhances robustness by randomly selecting top-ranked, low-ranked, or random passages during training.
+  • Finding: We demonstrate that the marginal benifit from sophisticated robust training diminish significantly as model capacity increases.
+  • Rationale: Through detailed analysis, we attribute this finding to intrinsic properties of stronger models-including better confidence calibration, cross-task generalization, and effective attention patterns—even under simple robust training.
+2. Preliminaries
+2.1. Related Work
+2.1.1. Retrieval-Augmented Generation
+Retrieval-Augmented Generation has been developed to enhance the reasoning and generation capabilities of LLMs by integrating external information.
+However, dense retrievers are not always perfect and can sometimes recall irrelevant or erroneous documents, which can degrade performance (Jiang et al., 2023; Ding et al., 2024).
+To mitigate this issue, adaptive retrieval techniques and robust RAG training methods have been proposed.
+Adaptive retrieval techniques retrieve external information only when LLMs encounter unknowns, reducing the risk of integrating misleading or incorrect information (Jiang et al., 2023; Mallen et al., 2023; Asai et al., 2023).
+Additionally, robust training focuses on enhancing LLMs’ response to various retrieval noises, aiming to efficiently obtain adversarial examples that improve model robustness while reducing training overhead (Fang et al., 2024; Zhu et al., 2024b; Xiang et al., 2024).
+2.1.2. Robust Training for RAG Systems
+The performance of LLMs in RAG systems can be compromised by irrelevant context (Cuconasu et al., 2024; Wu et al., 2024).
+Recently, researchers have focused on this issue and proposed effective adversarial training methods to enhance the robustness of LLMs against noisy documents.
+RetRobust (Yoran et al., 2024) suggests using a mix of relevant and irrelevant documents to train LLMs to withstand noisy contexts.
+RAAT (Fang et al., 2024) and ATM (Zhu et al., 2024b) employ adversarial training to dynamically adjust the model’s training process, aiming to maintain robustness against noise and generate accurate answers.
+RobustRAG (Xiang et al., 2024) proposes to enhances RAG model robustness against retrieval corruption attacks through an isolate-then-aggregate strategy to achieve certifiable robustness.
+Additionally, another work focus on identifying key documents from noisy contexts to facilitate effective question answering through knowledge rewriting, refinement, or filtering (Zhu et al., 2024a; Qiao et al., 2024; Jin et al., 2024; Zeng et al., 2024).
+However, these robust training approaches are primarily applied to small or weak LMs with fewer than 7 billion parameters.
+Thus, there’s an urgent need to explore whether complex robust training is still necessary to improve the robustness and generalization of bigger or stronger models when dealing with noisy contexts.
+2.2. Robustness Training for RAG
+2.2.1. Retrieval-augmented Adaptive Adversarial Training
+To improve the robustness of retrieval-augmented language models (RALMs) against retrieval noise, RAAT (Fang et al., 2024) incorporates an adversarial loss into the standard supervised fine-tuning (SFT).
+The model processes one golden context and three adversarial samples at each iteration, optimizing for the most challenging perturbation.
+The adversarial objective follows a min-max strategy:
+(1) $\mathcal{L}_{\text{max}}=\max_{da\in DA}\mathcal{L}(\theta,da(q),a),$
+where $\mathcal{L}$ represents the generation loss function to noisy retrievals, and $q^{\prime}=da(q)$ represents the augmented noise context of $q$.
+A regularization term controls excessive sensitivity to noise:
+(2) $\mathcal{L}_{\text{ada}}=\mathcal{L}_{\text{max}}+w_{\text{reg}}\cdot|\mathcal{L}_{\text{max}}-\mathcal{L}_{\text{min}}|_{2}^{2},$
+The adversarial loss $\mathcal{L}_{\text{ada}}$ combines both terms, where the regularization term helps stabilize training by mitigating excessive sensitivity to retrieval noise.
+The regularization term, calculated as the square of the difference between $\mathcal{L}_{\text{max}}$ and $\mathcal{L}_{\text{min}}$, encourages a more balanced optimization, preventing the model from overreacting to the most challenging perturbations.
+2.2.2. Multiagent Iterative Tuning Optimization
+ATM (Zhu et al., 2024b) steers the generator to have a robust perspective of useful documents for question answering with the help of an auxiliary attacker agent.
+The generator is trained to maximize answer correctness while minimizing its sensitivity to adversarial perturbations.
+It receives the user query along with a document list, which may contain both relevant and fabricated information introduced by the attacker.
+Its objective is to generate accurate responses as long as sufficient truthful information is present while reducing the impact of misleading or fabricated content.
+To achieve this, the generator learns to identify and leverage relevant documents while ignoring noisy ones, regardless of whether they originate from the original retrieval $\mathcal{D}$ or adversarial perturbations $\mathcal{D}^{\prime}$.
+This can be formalized as maximizing the objective:
+(3) $G(a\mid q,\mathcal{D}^{\prime})-\mathrm{dist}\left[G(a\mid q,\mathcal{D}),G(a\mid q,\mathcal{D}^{\prime})\right],$
+where ${G}(\cdot)$ represents the language model probability of generating an answer, and $\mathrm{dist}\left[\cdot\right]$ measures the divergence between outputs under different document conditions.
+The generator and the attacker are tuned adversarially for several iterations.
+After rounds of multi-agent iterative tuning, the generator can eventually better discriminate useful documents amongst fabrications.
+2.2.3. Invariant Risk Minimization
+Invariant Risk Minimization (IRM) aims to learn representations that remain stable across different environments, improving generalization under distribution shifts.
+To enhance the robustness of retrieval-augmented generation (RAG) across varying retrieval conditions, the V-REx objective (Krueger et al., 2021) can also be adapted to enforce risk invariance across different retrieval environments.
+Given retrieval environments $\mathcal{E}=\{1,\dots,m\}$, where each $e\in\mathcal{E}$ corresponds to a specific retrieval scenario (e.g., golden documents, top-$k$ retrieved, noisy retrieval), we define the empirical risk $\mathcal{R}_{e}(\theta)$ of the model parameterized by $\theta$.
+The training objective is formulated as:
+(4) $\mathcal{R}_{\textrm{V-REx-RAG}}(\theta)=$ $\beta\;\mathrm{Var}(\{\mathcal{R}_{1}(\theta),.,\mathcal{R}_{m}(\theta)\})$
+$+\sum^{m}_{e=1}\mathcal{R}_{e}(\theta),$
+where $\beta\geq 0$ controls the trade-off between minimizing average risk and enforcing risk invariance across retrieval environments.
+A higher $\beta$ reduces performance discrepancies caused by retrieval variations, improving generalization under both high-quality and noisy retrieval conditions.
+3. Evaluating Marginal Benefit of Robust RAG Methods
+In this section, we establish a systematic evaluation framework to address a fundamental question in RAG: whether sophisticated robust training methods continue to yield substantial performance improvements as model capabilities increase.
+We first formalize our evaluation metric to quantify the marginal benefit provided by different robust RAG strategies, followed by a detailed categorization of current state-of-the-art robust RAG methods.
+We also provide an experimental setup that ensures fair comparison across model scales and dataset types.
+3.1. Metric: Marginal Robustness Benefit
+To measure the relative utility of sophisticated robust training methods, we introduce the Marginal Robustness Benefit, denoted as $\Delta$, which captures the relative performance gain between the best and worst robust RAG strategies:
+(5) $\Delta=\frac{S_{\text{best}}-S_{\text{worst}}}{S_{\text{worst}}}$
+where $S_{\text{best}}$ and $S_{\text{worst}}$ denote the performance scores (EM or F1) of the most and least effective training strategies, respectively.
+By computing $\Delta$ across models of varying capacities, we quantify how the relative effectiveness of robust training evolves with model scaling.
+This metric helps assess whether the complexity introduced by advanced training strategies remains necessary as foundation models grow more capable.
+3.2. Taxonomy of Robust RAG Methods
+We categorize current robust RAG training approaches into two dimensions:
+3.2.1. Document Selection Strategy
+We first analyze robust RAG training methods based on document selection strategies.
+These approaches use different criteria to select documents for supervised fine-tuning, thereby improving model adaptability to retrieval variations:
+  • Golden Document: Selects the most relevant document containing the correct answer. If none exist in the top-20 documents, use the top-1 document for consistency.
   • Top-1 Document: Uses the highest-scoring retrieved document, which may not contain the correct answer, reflecting real-world retrieval challenges.
-  • Golden Document: Selects the most relevant document containing the correct answer. If none exist in the top-20, it defaults to the top-1 document for consistency.
-Model RAG Scenario HotpotQA NQ WebQuestions TriviaQA AVERAGE
+  • RALM: Fine-tunes the model by prepending golden retrieval text. Queries without golden documents in the top-20 docs are excluded.
+  • RetRobust: Following Yoran et al. (2024), this method enhances robustness by randomly selecting top-ranked, low-ranked, or random passages for training.
+To assess robustness against irrelevant information, we also include two adversarial scenarios:
+  • Random Document: Randomly selects a document from retrieved results.
+  • Irrelevant Document: Uses a passage from another query’s results, ensuring complete irrelevance.
+3.2.2. Adversarial Loss Design
+Next, we analyze robust RAG methods based on adversarial loss designs.
+These approaches enhance model resilience against retrieval noise by combining standard SFT with robustness-enforcing regularization.
+Models train across multiple retrieval environments (golden documents, top-k retrievals, or adversarially perturbed documents) using a dual objective: standard training loss plus regularization to reduce sensitivity to retrieval variations.
+Dataset Type # Train # Dev
+NQ single-hop 79,168 8,757
+WebQuestions single-hop 2,474 278
+TriviaQA multi-hop 78,785 8,837
+HotpotQA multi-hop 90,447 7,405
+Table 1. Statistics of different datasets
+We evaluate two adversarial loss strategies:
+  • RAAT: Reduces performance gap between best and worst retrieval cases, ensuring stability under challenging conditions.
+  • IRM: Minimizes performance variance across retrieval environments, mitigating sensitivity to distribution shifts for reliable cross-scenario performance.
+3.3. Experimental Setups
+Model RAG Methods HotpotQA NQ WebQuestions TriviaQA AVERAGE
 EM F1 EM F1 EM F1 EM F1 EM F1
-Llama-2-7b-chat-hf Base Model 3.30 12.34 1.21 10.61 0.00 13.08 4.32 20.27 2.21 14.08
+Llama-2-7b-chat-hf
+Base Model 3.30 12.34 1.21 10.61 0.00 13.08 4.32 20.27 2.21 14.08
 RALM 26.21 36.42 32.17 42.68 33.81 45.85 50.28 60.17 35.62 46.28
@@
 Golden Doc 30.67 42.78 36.50 47.77 39.93 52.11 50.25 63.28 39.34 51.49
-[…]
+Random Doc 30.94 43.11 38.16 49.78 42.45 53.97 52.72 65.52 41.07 53.10
 Irrelevant Doc 31.01 42.98 37.08 48.93 39.21 50.79 51.97 64.70 39.82 51.85
-[…]
+RAAT 31.32 43.24 42.91 53.19 36.69 48.82 51.65 58.71 40.64 50.99
 IRM 34.38 47.11 40.96 53.07 53.96 61.62 57.58 69.08 46.72 57.72
-$\Delta$ (Worst $\rightarrow$ Best) 16.19% 12.89% 21.87% 13.19% 48.53% 28.43% 13.55% 17.66% 15.95% 13.20%
-Llama-3-8B-Instruct Base Model 23.31 32.60 30.04 41.59 26.98 43.25 58.80 66.45 34.78 45.97
-RALM 27.64 38.10 35.19 46.10 47.84 56.98 54.75 63.27 41.36 51.11
-RetRobust 36.06 48.99 43.28 55.04 52.88 62.10 59.06 67.77 47.82 58.48
-Top-1 Doc 36.72 49.30 44.38 56.20 54.68 62.26 60.80 68.31 49.15 59.02
-Golden Doc 35.52 48.31 41.35 53.13 48.92 58.41 58.26 66.99 46.01 56.71
-[…]
-Irrelevant Doc 35.31 47.92 42.45 54.41 46.76 57.67 58.97 66.57 45.87 56.64
-[…]
-IRM 35.19 48.08 41.13 53.14 53.96 61.64 57.15 69.13 46.86 58.00
-$\Delta$ (Worst $\rightarrow$ Best) 14.04% 12.53% 7.90% 5.76% 16.94% 8.46% 11.74% 10.70% 10.92% 8.42%
-Table 1: Performance comparison of different LLMs (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) across different robust RAG scenarios on four datasets (HotpotQA, NQ, WebQuestions, and TriviaQA).
-The row $\Delta$ (Worst $\rightarrow$ Best) indicates the performance gain achieved by the best method compared to the worst strategy among these training strategies, representing the benefit of sophisticated robust training methods.
-To assess robustness against irrelevant information, we introduce two adversarial scenarios:
-  • Random Document: Randomly selects a document from retrieved results, simulating unpredictable retrieval quality.
-  • Irrelevant Document: Chooses a passage from another query’s retrieval results, ensuring no relevance to the current query.
-3.3.2 Adversarial Loss Design
-We assess two popular adversarial loss strategies:
-  • RAAT.
-The regularization term in RAAT reduces the performance gap between the best and worst retrieval cases.
-By penalizing excessive performance disparity, RAAT ensures that the model remains stable even under challenging retrieval conditions, leading to improved robustness and generalization.
-  • IRM.
-The regularization in IRM minimizes the variance in performance across different retrieval environments.
-By enforcing consistency, IRM mitigates sensitivity to distribution shifts, ensuring that the model performs reliably across diverse retrieval scenarios.
-4 Does Sophisticated Robust Training Still Matter in Powerful Models?
+$\Delta$ (Worst $\rightarrow$ Best) 31.17% 29.35% 33.39% 24.63% 59.60% 34.39% 14.65% 17.66% 31.16% 24.72%
+Llama-3-8B-Instruct
+Base Model 23.31 32.60 30.04 41.59 26.98 43.25 58.80 66.45 34.78 45.97
+RALM 30.73 41.25 35.19 46.10 47.84 56.98 54.75 63.27 41.36 51.11
+RetRobust 36.88 49.43 43.28 55.04 52.88 62.10 59.06 67.77 48.03 58.59
+Top-1 Doc 37.09 49.67 44.38 56.20 54.68 62.26 60.80 68.31 49.24 59.11
+Golden Doc 35.68 48.97 41.35 53.13 48.92 58.41 58.26 66.99 46.05 56.88
+Random Doc 36.73 49.53 43.37 55.43 53.24 62.55 60.62 68.64 48.49 59.04
+Irrelevant Doc 35.53 48.27 42.45 54.41 46.76 57.67 58.97 66.57 45.93 56.73
+RAAT 32.79 44.26 42.34 53.31 48.28 58.17 54.41 62.45 44.46 54.55
+IRM 35.63 48.72 41.13 53.14 53.96 61.64 57.15 69.13 46.97 58.16
+$\Delta$ (Worst $\rightarrow$ Best) 20.70% 20.41% 26.12% 21.91% 16.94% 9.78% 11.74% 10.70% 16.88% 13.89%
+Table 2.
+Performance comparison of Llama-2-7b-chat-hf and Llama-3-8B-Instruct across various robust RAG methods on four datasets (HotpotQA, NQ, WebQuestions, and TriviaQA).
+The row $\Delta$ (Worst $\rightarrow$ Best) indicates the marginal robustness benefit achieved by the best method compared to the worst method.
+Model RAG Methods HotpotQA NQ WebQuestions TriviaQA AVERAGE
+EM F1 EM F1 EM F1 EM F1 EM F1
+Qwen1.5-7B-Chat
+Base Model 24.15 34.06 24.31 35.27 23.02 38.65 48.43 57.65 29.98 41.41
+RALM 24.44 34.23 29.83 40.22 44.24 53.46 46.40 55.96 36.23 45.97
+RetRobust 29.05 40.99 34.82 45.92 47.12 55.39 49.16 59.01 40.04 50.33
+Top-1 Doc 29.68 41.57 34.16 45.44 43.88 52.71 49.61 59.03 39.33 49.69
+Golden Doc 28.66 40.55 32.93 43.99 43.88 52.74 49.25 58.87 38.68 49.04
+Random Doc 28.83 40.66 33.96 45.09 46.04 53.21 48.75 58.54 39.40 49.38
+Irrelevant Doc 28.12 39.87 32.74 43.94 43.88 52.67 48.33 58.18 38.27 48.67
+IRM 26.95 38.67 30.59 41.83 47.84 55.68 45.52 56.24 37.73 48.11
+$\Delta$ (Worst $\rightarrow$ Best) 21.44% 21.44% 16.73% 14.17% 9.02% 5.71% 8.99% 5.49% 10.52% 9.48%
+Qwen2.5-7B-Instruct
+Base Model 25.10 35.02 26.97 38.15 25.90 42.46 53.86 62.73 32.96 44.59
+RALM 26.37 36.30 31.43 42.00 41.73 51.98 53.92 62.74 38.36 48.26
+RetRobust 30.47 42.26 34.60 46.00 45.68 53.80 54.43 63.63 41.30 51.42
+Top-1 Doc 30.24 42.61 34.72 46.20 46.04 54.13 53.65 63.01 41.16 51.49
+Golden Doc 30.28 41.84 33.77 45.11 44.60 53.18 54.35 63.55 40.75 50.92
+Random Doc 30.25 42.22 34.08 45.47 44.24 53.30 54.29 63.54 40.72 51.13
+Irrelevant Doc 29.76 41.67 33.64 44.87 42.09 52.72 53.38 63.01 39.72 50.57
+IRM 29.58 41.73 32.79 44.17 46.40 54.80 53.45 64.26 40.56 51.24
+$\Delta$ (Worst $\rightarrow$ Best) 15.55% 17.38% 10.47% 10.00% 11.19% 5.43% 1.97% 2.42% 7.64% 6.70%
+Table 3. Performance comparison of different LLMs (Qwen1.5-7B-Chat and Qwen2.5-7B-Instruct) across different robust RAG scenarios on four datasets (HotpotQA, NQ, WebQuestions, and TriviaQA).
+Datasets
+For our experiment, we evaluate on four widely-used question answering datasets: (1) single-hop QA, including NaturalQuestions (NQ) (Kwiatkowski et al., 2019) and WebQuestions (Berant et al., 2013); and (2) multi-hop QA, including TriviaQA (Joshi et al., 2017) and HotpotQA (Yang et al., 2018).
+All experimental results are evaluated on their dev splits using the Exact Match (EM) and F1 metrics.
+Detailed statistics of these datasets are listed in Table 1.
+Implementation Details
+We implement a standard RAG framework with retrieval and generation phases.
+For retrieval, we use Contriever (Izacard et al., 2022), a BERT-based dense retriever to retrieve top-20 documents per query as a candidate pool.
+We sample different document subsets from the candidate pool to create varying retrieval environments for each methods.
+For question-answering, the top-5 retrieved documents are concatenated with the query as input for answer generation.
+For model training, we utilize the LLaMA-Factory library to facilitate efficient LLM finetuning.
+We train our models with a learning rate of 1e-6 and set the maximum number of epochs to 3.
+To optimize the training process, we employ bfloat16 precision and DeepSpeed ZeRO-3 for distributed training.
+The gradient accumulation steps are set to 2 with a batch size of 8.
+For inference, we leverage vLLM to ensure efficient model serving.
+We maintain the default decoding parameters for each model during inference, including the default temperature and top-p sampling probabilities, to ensure fair comparison across different model variants.
+All training and inference procedures are conducted on 8 $\times$ NVIDIA A100 80GB GPUs.
+4. Diminishing Benefits of Complex Methods With Model Capacity
 To investigate whether sophisticated document selection strategies and adversarial loss designs are still essential for robust RAG performance as LLMs continue to evolve, we conduct comprehensive experiments across multiple LMs and datasets.
-4.1 Do Sophisticated Document Selection Strategies Matter?
-We conduct experiments to analyze the effectiveness of complex document selection strategies under Llama model families (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) in Table 1, and Qwen model families in Appendix Table 4.
+4.1. Marginal Benefits of Sophisticated Document Selection
+We conduct experiments to analyze the effectiveness of complex document selection strategies under Llama model families (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) in Table 2 and Qwen model families (Qwen1.5-7B-Chat and Qwen2.5-7B-Instruct) in Table 3.
 Training with sophisticated documents enhances LM robustness for weak models
-The experimental results presented in Tables 1 and 4 provide compelling evidence that robust training significantly improves model resilience when processing noisy documents.
+The experimental results provide compelling evidence that robust training significantly improves model resilience when processing noisy documents.
 While base models exhibit substantial performance degradation when encountering noisy documents during inference, models that undergo robust training maintain consistent and superior QA performance across various document selection strategies.
@@
 Diminishing returns of sophisticated document selections as models evolve
-Experimental results indicate that the performance gains from sophisticated document selection strategies diminish as models evolve.
-For instance, comparing Llama-2-7b-chat-hf with Llama-3-8B-Instruct, the improvement in performance due to advanced document selection strategies decreases significantly, with the $\Delta$ (Worst $\rightarrow$ Best) metric for NQ dropping from 21.87% to 7.90% EM.
-A similar trend is observed when comparing Qwen1.5-7B-Chat to Qwen2.5-7B-Instruct, where the performance improvement from sophisticated document selection also shows a noticeable reduction.
+Experimental results indicate that the marginal robust benefit from sophisticated document selection strategies diminish as models evolve.
+For instance, comparing Llama-2 with Llama-3, the improvement in performance due to advanced document selection strategies decreases significantly, with the $\Delta$ (Worst $\rightarrow$ Best) metric for WebQuestions dropping from 59.60% to 16.94% EM.
+A similar trend is observed when comparing Qwen1.5 to Qwen2.5, where the performance improvement from sophisticated document selection also shows a noticeable reduction.
 These results suggest that as models become more advanced, their ability to process and utilize information improves independently of complex document selection strategies, leading to diminished returns from such methods.
-4.2 Do Adversarial Loss Functions Matter?
-To investigate whether the design of complex adversarial loss functions contributes to model performance, in Table 1 and 4, we also analyze the robustness of various adversarial loss designs.
+4.2. Marginal Benefits of Complex Loss Design
+To investigate whether the design of complex adversarial loss functions contributes to model performance, in Table 2 and 3, we also analyze the robustness of various adversarial loss designs.
 Adversarial loss significantly enhances performance for weaker models
@@
 Notably, in some cases, RAAT and IRM outperform traditional document selection strategies (e.g., top-1 Doc and golden doc), demonstrating their value in scenarios where the model needs stronger guidance to handle noisy retrievals.
-Figure 2: Comparison of F1 Scores on the HotpotQA dataset for training with Random Doc and Golden Doc across models with varying parameter sizes from 0.5B to 70B.
-We provide detailed results in Appendix Table 6.
+Figure 2. Comparison of F1 Scores on the HotpotQA (left) and NQ (right) dataset for training with Random Doc and Golden Doc across models with varying parameter sizes from 0.5B to 70B.
 Adversarial loss exhibits diminishing returns for stronger models
 For the stronger models, such as Llama-3-8B-Instruct and Qwen2.5-7B-Instruct, the benefits of adversarial loss functions are less pronounced.
-The basic Llama-3-8B-Instruct already achieves an average EM / F1 of 34.78 / 45.97, and the introduction of adversarial losses (RAAT and IRM) results in only modest improvements, with average EM / F1 scores of 44.31 / 54.44 and 46.86 / 58.00, respectively.
-Similarly, Qwen2.5-7B-Instruct shows similar trends, with marginal gains from adversarial loss functions.
-These improvements are comparable to or even slightly worse than the performance achieved by random document selection (48.30 / 58.92) or top-1 document strategies (49.15 / 59.02).
+The basic Llama-3-8B-Instruct already achieves an average EM / F1 of 34.78 / 45.97, and the introduction of adversarial losses (RAAT and IRM) results in only modest improvements, with average EM / F1 scores of 44.46 / 54.55 and 46.97 / 58.16, respectively.
 This indicates that the models’ inherent robustness reduces the impact of adversarial losses, and in some cases, may even hinder performance.
 We hypothesize that when the model’s internal robustness is already well-developed, additional constraints from adversarial losses may interfere with its ability to optimize on clean and relevant inputs.
+Conclusion: the necessity of adversarial loss.
 Based on these findings, we conclude that both adversarial loss functions and document selection strategies are more beneficial for weaker models.
@@
 This underscores the importance of tailoring training strategies to the model’s inherent capabilities.
-4.3 Do Training Strategies Matter Across Model Scales?
-Through analyzing Tables 1 and 4, we observe a counter-intuitive phenomenon: models trained with random documents outperformed those with golden documents, despite the latter containing ground truth answers typically yielding optimal results in standard SFT.
+4.3. Scaling Behavior and Cross-Domain Generalization
+In this section, we conduct experiments across a wide range of model scales and diverse domains to investigate how the effectiveness of robust RAG training strategies varies with model capabilities.
+4.3.1. Parameter Scale Analysis
+Through analyzing Tables 2 and 3, we observe a counter-intuitive phenomenon: models trained with random documents outperformed those with golden documents, despite the latter containing ground truth answers typically yielding optimal results in standard SFT.
 To investigate whether this phenomenon extends beyond 7B-8B models, we conduct experiments across model scales from 0.5B to 70B parameters.
-The results (detailed in Appendix Table 6) in Figure 2 demonstrate that for smaller models ($\leq$3B parameters), training with golden documents leads to superior performance.
+The results in Figure 2 demonstrate that for smaller models ($\leq$3B parameters), training with golden documents leads to superior performance.
 This suggests that smaller models, limited by their inherent capabilities, benefit more from high-quality golden documents containing direct answers.
@@
 This finding indicates that sophisticated document selection strategies become less crucial as model size increases, revealing an important scaling property in model training.
-5 Why Sophisticated Training No Longer Matters in Powerful Models?
+4.3.2. Cross-Domain Generalization
+To validate our findings beyond general question-answering tasks, we examine whether this phenomenon generalizes across temporal reasoning (TimeQA (Chen et al., 2021)), specialized legal knowledge (LegalBench (Guha et al., 2023)), and scenarios with conflicting evidence (RAGuard (Zeng et al., 2025)).
+Model Method TimeQA
+EM F1
+Llama-2-7b-chat-hf Golden Doc 32.93 43.44
+Top-1 Doc 32.26 43.13
+Random Doc 31.90 43.32
+RetRobust 31.97 42.58
+$\Delta$ (Worst $\rightarrow$ Best) 3.23% 2.02%
+Llama-3-8B-Instruct Golden Doc 35.44 46.90
+Top-1 Doc 35.27 46.94
+Random Doc 35.87 47.20
+RetRobust 35.90 47.15
+$\Delta$ (Worst $\rightarrow$ Best) 1.79% 0.63%
+Table 4. Performance comparison of different LLMs (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) across various robust RAG methods on the TimeQA dataset.
+Temporal Reasoning Domain
+We evaluate the performance of various document selection strategies on the TimeQA dataset (Chen et al., 2021), which specifically tests models’ ability to handle time-sensitive information.
+As shown in Table 4, we compare two different language models: Llama-2-7b-chat-hf and Llama-3-8B-Instruct.
+With Llama-2-7b-chat-hf, we observe moderate variations across strategies, with a delta of 3.23% in EM and 2.02% in F1 between best and worst approaches.
+Golden Document selection generally yields the strongest performance.
+However, with Llama-3-8B-Instruct, this marginal robustness benefit narrows significantly to just 1.79% for EM and 0.63% for F1, with RetRobust and Random Document selection slightly outperforming Golden Document selection.
+Model Method LegalBench
+Llama-2-7b-chat-hf Base Model 2.55
+RALM 70.96
+RetRobust 66.67
+Top-1 Doc 71.21
+Golden Doc 70.96
+Random Doc 70.96
+Irrelevant Doc 72.73
+IRM 69.19
+$\Delta$ (Worst $\rightarrow$ Best) 9.09%
+Llama-3-8B-Instruct Base Model 73.23
+RALM 85.61
+RetRobust 84.85
+Top-1 Doc 85.86
+Golden Doc 86.62
+Random Doc 86.87
+Irrelevant Doc 86.36
+IRM 83.84
+$\Delta$ (Worst $\rightarrow$ Best) 3.61%
+Table 5. Performance comparison of the Llama-2-7b-chat-hf and Llama-3-8B-Instruct models across various robust RAG methods on the Legal-Bench Consumer Contracts QA dataset.
+Specialized Legal Domain
+We extend our investigation to specialized domains by evaluating various robust RAG methods on the LegalBench Consumer Contracts QA dataset (Guha et al., 2023).
+Table 5 presents a comprehensive comparison across both Llama-2-7b-chat-hf and Llama-3-8B-Instruct models.
+The results strongly reinforce our core findings.
+For Llama-2-7b-chat-hf, we observe substantial performance variations across different methods, with a 9.09% delta between the best (Irrelevant Doc at 72.73%) and worst (RetRobust at 66.67%) effective approaches.
+This significant variation suggests that older models remain highly sensitive to the specific training methodology employed.
+In contrast, the more advanced Llama-3-8B-Instruct model demonstrates higher baseline performance (73.23% without retrieval augmentation) and reduced sensitivity to the choice of method.
+The performance delta between the best (Random Doc at 86.87%) and worst effective method (IRM at 83.84%) narrows to just 3.61% - less than half the gap observed with Llama-2.
+Model Method RAGuard
+Llama-2-7b-chat-hf
+Random Doc 56.04
+Top-1 Doc 56.73
+Golden Doc 54.31
+Irrelevant Doc 56.00
+IRM 51.81
+RetRobust 46.71
+$\Delta$ (Worst $\rightarrow$ Best) 21.45%
+Llama-3-8B-Instruct
+Random Doc 59.82
+Top-1 Doc 60.08
+Golden Doc 59.37
+Irrelevant Doc 59.14
+IRM 58.95
+RetRobust 59.78
+$\Delta$ (Worst $\rightarrow$ Best) 1.92%
+Table 6. Performance comparison of different LLMs (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) across various robust RAG methods on the RAGuard Benchmark.
+Evaluation with Conflicting Evidence
+To validate our findings in challenging scenarios, we evaluate robust RAG methods on the RAGuard benchmark (Zeng et al., 2025), which tests model robustness against contradictory retrievals.
+Table 6 shows that for Llama-2-7b-chat-hf, there is a substantial 21.45% performance gap between the best and worst training methods.
+In contrast, Llama-3-8B-Instruct demonstrates remarkably consistent performance, with the gap shrinking to merely 1.92% - a 90% reduction compared to the less capable model.
+These results on a benchmark with conflicting evidence further confirm our central finding that the benefits of complex robust training methods diminish significantly as model capacity increases.
+Across all three domains, we consistently observe that performance variations between different document selection strategies decreases significantly with more advanced models.
+These experiments confirm that as language models increase in capability, the benefits of complex robust RAG training strategies diminish significantly, regardless of domain specificity.
+5. Understanding Diminishing Benefits of Complex Training in Powerful LLMs
 In this section, we conduct comprehensive experiments to delve into the reasons why sophisticated robust training strategies may no longer be crucial in powerful models.
-Figure 3: Confidence scores for correct and wrong answers on HotpotQA dataset, comparing Llama2 and Llama3 models across various robust training methods.
-5.1 Powerful Models Enable Natural Calibration
+Figure 3. Confidence scores for correct and wrong answers on HotpotQA dataset, comparing Llama2 and Llama3 models across various robust training methods.
+5.1. Powerful Models Enable Natural Calibration
 To understand whether powerful models inherently possess the ability to distinguish reliable from unreliable answers, we take HotpotQA dataset as an example to examine the confidence calibration capabilities for Llama2 (Llama-2-7b-chat-hf) and Llama3 (Llama-3-8B-Instruct) models.
-Here, confidence is the mean of token-wise probabilities in the model’s generated answer, providing a measure of the model’s certainty in its predictions Varshney et al.
-(2023); Xiong et al.
-(2024).
+Here, confidence is the mean of token-wise probabilities in the model’s generated answer, providing a measure of the model’s certainty in its predictions (Varshney et al., 2023; Xiong et al., 2024).
 Figure 3 reveals striking differences in their calibration patterns.
@@
 In contrast, Llama3 demonstrates inherently better calibration, maintaining higher confidence for correct answers (97.5) than incorrect ones (91.3) without any specialized complex training.
-While robust training methods (Golden Doc, Random Doc, and IRM) can effectively calibrate confidence scores and improve the gap between correct and incorrect answers for Llama2 from -2 to 12 with IRM, the marginal benefits of these complex training strategies diminish as model architectures advance.
+While robust training methods can effectively calibrate confidence scores and improve the gap between correct and incorrect answers for Llama2 from -2 to 12 with IRM, the marginal benefits of these complex training strategies diminish as model architectures advance.
 For Llama3, which already achieves a 6.2 confidence gap naturally, the improvements from these training methods become less significant.
 This finding strongly suggests that advances in model architecture can effectively eliminate the need for complex robustness training procedures, as newer models come with better built-in calibration capabilities.
-Figure 4: Generalization performance comparison across different strategies trained on HotpotQA (diagonal hatches bars) and evaluated on NQ, WebQuestions, TriviaQA datasets (plain bars).
-More results available in Appendix Table 8.
-5.2 Simple Training Strategies Generalize Well in Powerful Models
+Figure 4. Generalization performance comparison across different strategies trained on HotpotQA (diagonal hatches bars) and evaluated on NQ, WebQuestions, TriviaQA datasets (plain bars).
+5.2. Simple Training Strategies Generalize Well in Powerful Models
 We further investigate whether powerful models can maintain robust generalization across different datasets with simple training strategies.
 We fine-tune models on HotpotQA using four document selection approaches and evaluate their transfer performance on NQ, WebQuestions, and TriviaQA.
+Figure 5.
+Attention visualization for a QA case.
+Each subplot shows attention distribution heatmaps across different models, where cell (ii, $j$) represents the average attention weight from the $j$-th attention layer to document ii.
+Text highlighted in green indicates the correct answer and corresponding Doc1, blue indicates key terms from the query, and red indicates incorrect model predictions.
+The color intensity in the heatmaps indicates attention strength.
 As shown in Figure 4, simple strategies demonstrate surprisingly strong generalization ability.
-Random document selection matches or even outperforms sophisticated IRM across all evaluation datasets, with performance gaps of less than 1%.
+Random document selection matches or even outperforms sophisticated IRM across all evaluation datasets, with marginal robustness benefit of less than 1%.
 For instance, in TriviaQA, random selection (69.5 F1) slightly surpasses both golden (68.2 F1) and IRM (68.7 F1) approaches.
@@
 These findings provide strong evidence that as models become more powerful, sophisticated training strategies become increasingly unnecessary.
-Figure 5: Attention visualization for a QA case.
-Each subplot shows attention distribution heatmaps across different models, where cell ($i$, $j$) represents the average attention weight from the $j$-th attention layer to document $i$.
-Text highlighted in green indicates the correct answer and corresponding Doc1, blue indicates key terms from the query, and red indicates incorrect model predictions.
-The color intensity in the heatmaps indicates attention strength.
-5.3 Powerful Models Learn Effective Attention Patterns with Simple Training
+5.3. Powerful Models Learn Effective Attention Patterns with Simple Training
 To provide a direct understanding of why simple training can achieve good performance, we visualize attention distributions across different training strategies.
@@
 This finding provides direct evidence that powerful models can learn optimal attention mechanisms even with simple training strategies, making sophisticated training methods unnecessary.
-Figure 6: Performance comparison with different numbers of random documents during training. Increasing the number of random documents consistently improves model performance.
-Figure 7: Training curves comparison between random and golden document strategies using Llama-2-7b-chat-hf and LLama-3-8B-Instruct.
-5.4 Training with Random Docs: Better Performance and Faster Convergence
+Figure 6. Performance comparison with different numbers of random documents during training. Increasing the number of random documents consistently improves model performance.
+Figure 7. Training curves comparison between random and golden document strategies using Llama-2-7b-chat-hf and LLama-3-8B-Instruct.
+5.4. Training with Random Docs: Better Performance and Faster Convergence
 We investigate why random training proves good performance from two aspects.
@@
 This faster convergence with better performance holds true for both model scales, indicating that simpler random training actually enables more efficient learning in powerful language models.
-6 Insights and Future Directions
-Our findings reveal a fundamental shift in RAG system: as models become more powerful, the marginal benefits of sophisticated training strategies diminish significantly.
-This observation has several important insights:
-Simplified RAG architecture design
-For powerful models, simple retrieval strategies (even random selection) can achieve comparable performance to sophisticated approaches.
-This enables streamlined RAG architectures by replacing elaborate document filtering mechanisms with simpler retrieval methods, substantially reducing system complexity without sacrificing performance.
-Scalable RAG for open-domain tasks.
-Larger models demonstrate robustness against noisy retrieval, suggesting that open-domain RAG systems can function effectively with minimal retrieval supervision.
-Instead of enforcing strict filtering of retrieved documents, future large-scale RAG systems can leverage weakly supervised learning, incorporating large-scale web data and noisy retrieval results to improve generalization.
-Theoretical implications for model scaling.
-This research reveals a previously unexplored aspect of scaling laws: the diminishing returns of complex training strategies as models grow larger.
-This challenges current theoretical frameworks and calls for new ones that better explain how training requirements evolve with model scale.
-Broader impact on machine learning.
-Our findings suggest that as models become more powerful, practitioners should prioritize architectural improvements and data quality over complex training strategies.
-This insight could lead to more efficient resource allocation in model development across various applications, from computer vision to natural language processing.
-7 Conclusions
-In this study, we systematically investigate whether complex robust training strategies remain necessary for RAG systems as model capacity grows.
-Our extensive experiments consistently show that while sophisticated training methods significantly enhance weaker models’ performance, their benefits diminish dramatically in more powerful models.
-Through the systematic analysis, we find that advanced models inherently possess strong confidence calibration, cross-dataset generalization, and effective attention patterns even with simple training.
-These findings suggest that RAG systems can benefit from simpler training strategies as models become more powerful, enabling more scalable applications with minimal complexity.
-Limitations
-While our study provides valuable insights into the impact of adversarial loss functions and document selection strategies on RAG model robustness, several limitations remain.
-First, our analysis is restricted to dense transformer-based models, leaving the effectiveness of these techniques on sparse models, such as mixture-of-experts (MoE) architectures, unexplored.
-Future work could investigate whether similar trends hold for sparsely activated models with dynamic routing mechanisms.
-Second, although we analyze the effectiveness of adversarial training, we do not explicitly examine its long-term stability or convergence properties, which may vary depending on hyperparameter choices and optimization dynamics.
-Additionally, while we demonstrate that stronger models exhibit diminishing returns from adversarial losses and document selection strategies, the precise mechanisms behind this phenomenon remain unclear.
-Further research is needed to understand how model capacity interacts with retrieval robustness.
-Ethics Statements
-Our study focuses on improving the robustness of RAG models, but several ethical considerations must be acknowledged.
-First, while adversarial training enhances model reliability, it does not eliminate the risk of biased or misleading outputs, particularly when retrieval sources contain inherent biases or misinformation.
-Future work should explore fairness-aware adversarial training to mitigate potential harms.
-Second, our findings suggest that stronger models require less intervention in document selection and loss design, which may influence resource allocation in real-world applications.
-Researchers and practitioners should ensure that model improvements do not disproportionately benefit well-resourced institutions while leaving smaller models less robust.
-Lastly, our experiments are conducted on widely used benchmark datasets, which may not fully reflect the diversity of real-world information needs.
-We encourage further research on robustness evaluation across varied domains, including low-resource languages and specialized knowledge fields, to ensure equitable advancements in RAG technology.
-Appendix A Related Work
-A.1 Retrieval-Augmented Generation
-Retrieval-Augmented Generation (RAG) has been developed to enhance the reasoning and generation capabilities of LLMs by integrating external information.
-RAG improves the overall performance of LLMs by retrieving relevant documents that provide additional context and support for the generation process.
-However, dense retrievers are not always perfect and can sometimes recall irrelevant or erroneous documents, which can degrade performance Jiang et al.
-(2023); Deng et al.
-(2023); Ding et al.
-(2024).
-To mitigate this issue, adaptive retrieval techniques and robust RAG training methods have been proposed.
-Adaptive retrieval techniques retrieve external information only when LLMs encounter unknowns, reducing the risk of integrating misleading or incorrect information Jiang et al.
-(2023); Mallen et al.
-(2023); Asai et al.
-(2023).
-Additionally, robust training focuses on enhancing LLMs’ response to various retrieval noises, aiming to efficiently obtain adversarial examples that improve model robustness while reducing training overhead Fang et al.
-(2024); Zhu et al.
-(2024a); Xiang et al.
-(2024).
-A.2 Robust Training for RAG Systems
-The performance of LLMs in RAG systems can be compromised by irrelevant context Cuconasu et al.
-(2024); Wu et al.
-(2024).
-Recently, researchers have focused on this issue and proposed effective adversarial training methods to enhance the robustness of LLMs against noisy documents.
-RetRobust Yoran et al.
-(2024) suggests using a mix of relevant and irrelevant documents to train LLMs to withstand noisy contexts.
-RAAT Fang et al.
-(2024) and ATM Zhu et al.
-(2024a) employ adversarial training to dynamically adjust the model’s training process, aiming to maintain robustness against noise and generate accurate answers.
-RobustRAG Xiang et al.
-(2024) proposes to enhances RAG model robustness against retrieval corruption attacks through an isolate-then-aggregate strategy to achieve certifiable robustness.
-Additionally, there is a body of work focused on identifying key documents from noisy contexts to facilitate effective question answering through knowledge rewriting, refinement, or filtering Zhu et al.
-(2024b); Qiao et al.
-(2024); Jin et al.
-(2024); Zeng et al.
-(2024).
-However, these robust training approaches are primarily applied to small or weak LMs with fewer than 7 billion parameters.
-Thus, there’s an urgent need to explore whether complex robust training is still necessary to improve the robustness and generalization of bigger or stronger models when dealing with noisy contexts.
-Appendix B Robustness Training for RAG
-B.1 Retrieval-augmented Adaptive Adversarial Training
-To improve the robustness of retrieval-augmented language models (RALMs) against retrieval noise, RAAT Fang et al.
-(2024) incorporates an adversarial loss into the standard supervised fine-tuning (SFT).
-The model processes one golden context and three adversarial samples at each iteration, optimizing for the most challenging perturbation.
-The adversarial objective follows a min-max strategy:
-$\mathcal{L}_{\text{max}}=\max_{da\in DA}\mathcal{L}(\theta,da(q),a),$ (2)
-where $\mathcal{L}$ represents the generation loss function to noisy retrievals, and $q^{\prime}=da(q)$ represents the augmented noise context of $q$.
-A regularization term controls excessive sensitivity to noise:
-$\mathcal{L}_{\text{ada}}=\mathcal{L}_{\text{max}}+w_{\text{reg}}\cdot|\mathcal{L}_{\text{max}}-\mathcal{L}_{\text{min}}|_{2}^{2},$ (3)
-The adversarial loss $\mathcal{L}_{\text{ada}}$ combines both terms, where the regularization term helps stabilize training by mitigating excessive sensitivity to retrieval noise.
-The regularization term, calculated as the square of the difference between $\mathcal{L}_{\text{max}}$ and $\mathcal{L}_{\text{min}}$, encourages a more balanced optimization, preventing the model from overreacting to the most challenging perturbations.
-B.2 Multiagent Iterative Tuning Optimization
-ATM Zhu et al.
-(2024a) steers the generator to have a robust perspective of useful documents for question answering with the help of an auxiliary attacker agent.
-The generator is trained to maximize answer correctness while minimizing its sensitivity to adversarial perturbations.
-It receives the user query along with a document list, which may contain both relevant and fabricated information introduced by the attacker.
-Its objective is to generate accurate responses as long as sufficient truthful information is present while reducing the impact of misleading or fabricated content.
-To achieve this, the generator learns to identify and leverage relevant documents while ignoring noisy ones, regardless of whether they originate from the original retrieval $\mathcal{D}$ or adversarial perturbations $\mathcal{D}^{\prime}$.
-This can be formalized as maximizing the objective:
-$G(a\mid q,\mathcal{D}^{\prime})-\mathrm{dist}\left[G(a\mid q,\mathcal{D}),G(a\mid q,\mathcal{D}^{\prime})\right],$ (4)
-where ${G}(\cdot)$ represents the language model probability of generating an answer, and $\mathrm{dist}\left[\cdot\right]$ measures the divergence between outputs under different document conditions.
-The generator and the attacker are tuned adversarially for several iterations.
-After rounds of multi-agent iterative tuning, the generator can eventually better discriminate useful documents amongst fabrications.
-B.3 Invariant Risk Minimization
-Invariant Risk Minimization (IRM) aims to learn representations that remain stable across different environments, improving generalization under distribution shifts.
-To enhance the robustness of retrieval-augmented generation (RAG) across varying retrieval conditions, the V-REx objective Krueger et al.
-(2021) can also be adapted to enforce risk invariance across different retrieval environments.
-Given retrieval environments =$\mathcal{E}=\{1,\dots,m\}$, where each $e\in\mathcal{E}$ corresponds to a specific retrieval scenario (e.g., golden documents, top-$k$ retrieved, noisy retrieval), we define the empirical risk $\mathcal{R}_{e}(\theta)$ of the model parameterized by $\theta$.
-The training objective is formulated as:
-$\mathcal{R}_{\textrm{V-REx-RAG}}(\theta)=$ $\beta\;\mathrm{Var}(\{\mathcal{R}_{1}(\theta),...,\mathcal{R}_{m}(\theta)\})$ (5)
-$+\sum^{m}_{e=1}\mathcal{R}_{e}(\theta),$
-where $\beta\geq 0$ controls the trade-off between minimizing average risk and enforcing risk invariance across retrieval environments.
-A higher $\beta$ reduces performance discrepancies caused by retrieval variations, improving generalization under both high-quality and noisy retrieval conditions.
-Appendix C Appendix
-C.1 Dataset Statistics
-Detailed statistics of these datasets we used are listed in Table 2.
-Dataset Type # Train # Dev
-NQ single-hop 79,168 8,757
-WebQuestions single-hop 2,474 278
-TriviaQA multi-hop 78,785 8,837
-HotpotQA multi-hop 90,447 7,405
-Table 2: Statistics of different datasets
-C.2 Implementation Details
-For model training, we utilize the LLaMA-Factory library to facilitate efficient LLM finetuning.
-We train our models with a learning rate of 1e-6 and set the maximum number of epochs to 3.
-To optimize the training process, we employ bfloat16 precision and DeepSpeed ZeRO-3 for distributed training.
-The gradient accumulation steps are set to 2 with a batch size of 8.
-For inference, we leverage vLLM to ensure efficient model serving.
-We maintain the default decoding parameters for each model during inference, including the default temperature and top-p sampling probabilities, to ensure fair comparison across different model variants.
-All training and inference procedures are conducted on 8x NVIDIA A100 80GB GPUs.
-C.3 Prompt Instruction
-We present the prompt used to guide LM inference on noisy contexts in Table 3.
-Specifically, we concatenate the top-5 retrieved documents into a single passage, append the input question afterward, and prompt the LM to generate a concise answer.
-You need to answer my question and complete the question-and-answer pair following the format provided in the example.
-The answers should be short phrases or entities, not full sentences.
-Here are some examples to guide you.
-Example 1:
-Question: What is the capital of France?
-Answer: Paris
-Example 2:
-Question: Who invented the telephone?
-Answer: Alexander Graham Bell
-Example 3:
-Question: Which element has the atomic number 1?
-Answer: Hydrogen
-### Retrieved Documents: {}
-### Question: {}
-### Answer:
-Table 3: The generator prompt we used in our experiments.
-Model RAG Scenario HotpotQA NQ WebQuestions TriviaQA AVERAGE
-EM F1 EM F1 EM F1 EM F1 EM F1
-Qwen1.5-7B-Chat Base Model 24.15 34.06 24.31 35.27 23.02 38.65 48.43 57.65 29.98 41.41
-RALM 24.44 34.23 29.83 40.22 44.24 53.46 46.40 55.96 36.23 45.97
-RetRobust 29.05 40.99 34.82 45.92 47.12 55.39 49.16 59.01 40.04 50.33
-Top-1 Doc 29.68 41.57 34.16 45.44 43.88 52.71 49.61 59.03 39.33 49.69
-Golden Doc 28.66 40.55 32.93 43.99 43.88 52.74 49.25 58.87 38.68 49.04
-[…]
-Irrelevant Doc 28.12 39.87 32.74 43.94 43.88 52.67 48.33 58.18 38.27 48.67
-[…]
-$\Delta$ (Worst $\rightarrow$ Best) 21.44% 21.44% 16.73% 14.17% 9.02% 5.71% 8.99% 5.49% 10.52% 9.48%
-Qwen2.5-7B-Instruct Base Model 25.10 35.02 26.97 38.15 25.90 42.46 53.86 62.73 32.96 44.59
-RALM 26.37 36.30 31.43 42.00 41.73 51.98 53.92 62.74 38.36 48.26
-RetRobust 30.47 42.26 34.60 46.00 45.68 53.80 54.43 63.63 41.30 51.42
-Top-1 Doc 30.24 42.61 34.72 46.20 46.04 54.13 53.65 63.01 41.16 51.49
-Golden Doc 30.28 41.84 33.77 45.11 44.60 53.18 54.35 63.55 40.75 50.92
-[…]
-Irrelevant Doc 29.76 41.67 33.64 44.87 42.09 52.72 53.38 63.01 39.72 50.57
-[…]
-$\Delta$ (Worst $\rightarrow$ Best) 15.55% 17.38% 10.47% 10.00% 11.19% 5.43% 1.97% 2.42% 7.64% 6.70%
-Table 4: Performance comparison of different LLMs (Qwen1.5-7B-Chat and Qwen2.5-7B-Instruct) across different robust RAG scenarios on four datasets (HotpotQA, NQ, WebQuestions, and TriviaQA).
-The row $\Delta$ (Worst $\rightarrow$ Best) represents the performance gain achieved by the best method compared to the worst strategy among these training strategies, highlighting the benefit of sophisticated robust training methods.
-Model Document Distribution HotpotQA NQ WebQuestion TriviaQA AVERAGE
-EM F1 EM F1 EM F1 EM F1 EM F1
-Llama-2-7b-chat-hf Base Model 3.30 12.34 1.21 10.61 0.00 13.08 4.32 20.27 2.21 14.08
-1 Golden Doc 30.67 42.78 36.50 47.77 39.93 52.11 50.25 63.28 39.34 51.49
-1 Random Doc 30.94 43.11 38.16 49.78 42.45 53.97 52.72 65.52 41.07 53.10
-0 Random + 3 Golden 31.79 44.03 36.68 47.90 43.02 53.43 56.55 64.97 42.01 52.58
-1 Random + 2 Golden 32.82 44.77 36.70 48.21 43.38 54.34 56.93 65.67 42.46 53.25
-2 Random + 1 Golden 33.32 45.45 37.07 48.67 44.10 54.82 57.56 65.90 43.01 53.71
-3 Random + 0 Golden 34.74 46.79 37.35 49.34 45.18 55.70 57.77 66.40 43.76 54.56
-Llama-3-8B-Instruct Base Model 23.31 32.60 30.04 41.59 26.98 43.25 58.80 66.45 34.78 45.97
-1 Golden Doc 35.52 48.31 41.35 53.13 48.92 58.41 58.26 66.99 46.01 56.71
-1 Random Doc 35.98 49.05 43.37 55.43 53.24 62.55 60.62 68.64 48.30 58.92
-0 Random + 3 Golden 35.96 49.01 40.88 52.91 50.25 59.13 54.95 66.48 45.51 56.88
-1 Random + 2 Golden 36.22 49.07 41.22 53.13 50.41 59.97 55.27 66.77 45.78 57.24
-2 Random + 1 Golden 36.77 49.59 41.75 53.48 51.49 61.02 55.39 66.74 46.35 57.71
-3 Random + 0 Golden 37.33 50.30 42.45 54.32 52.05 61.16 56.20 67.83 47.01 58.40
-Table 5: Experimental comparison of RAG robust training using different numbers of random documents and golden documents mixed together, with Llama-2-7b-chat-hf and Llama-3-8B-Instruct models across four datasets.
-Model RAG Scenario HotpotQA NQ
-EM F1 EM F1
-Qwen2.5-0.5B-Instruct RetRobust 17.69 26.63 17.16 25.44
-Top-1 Doc 18.28 27.38 18.00 26.52
-Golden Doc 19.11 28.04 19.32 27.49
-Random Doc 18.10 27.23 16.27 24.62
-Qwen2.5-1.5B-Instruct RetRobust 23.35 34.13 25.01 35.35
-Top-1 Doc 24.08 34.31 25.53 35.75
-Golden Doc 23.70 34.03 25.61 35.52
-Random Doc 23.50 33.68 24.63 34.89
-Qwen2.5-3B-Instruct RetRobust 26.29 37.34 29.46 40.38
-Top-1 Doc 26.17 37.33 29.86 40.67
-Golden Doc 26.17 37.16 30.08 40.41
-Random Doc 26.28 37.26 29.34 40.09
-Mistral-7B-Instruct-v0.2 RetRobust 35.92 49.03 43.19 55.51
-Top-1 Doc 35.80 49.08 44.35 56.29
-Golden Doc 34.67 47.46 41.52 53.41
-Random Doc 35.52 48.71 43.87 55.84
-Qwen2.5-32B-Instruct RetRobust 33.96 46.28 35.67 48.77
-Top-1 Doc 34.47 46.88 35.90 48.80
-Golden Doc 33.68 45.62 35.29 47.69
-Random Doc 33.81 46.31 36.09 49.03
-Llama-2-70b-chat-hf RetRobust 36.22 49.58 41.49 54.73
-Top-1 Doc 35.45 48.75 40.39 53.92
-Golden Doc 34.27 47.72 39.23 51.69
-Random Doc 34.90 48.31 40.39 53.92
-Llama-3-70B-Instruct RetRobust 40.78 54.67 45.84 59.25
-Top-1 Doc 40.61 55.37 46.05 59.61
-Golden Doc 40.65 54.15 44.80 57.85
-Random Doc 39.91 54.71 43.71 57.86
-Qwen2.5-72B-Instruct RetRobust 37.06 49.96 39.99 52.70
-Top-1 Doc 37.52 50.30 39.87 52.81
-Golden Doc 36.73 49.66 38.86 51.73
-Random Doc 36.75 50.05 39.95 52.86
-Table 6: Performance comparison of LLM with different parameter sizes (from 0.5B to 72B) in different robust RAG scenarios on HotpotQA and NQ datasets.
-Models RAG Scenario HotpotQA NQ WebQ TriviaQA
-Correct Wrong Correct Wrong Correct Wrong Correct Wrong
-Llama2-7b-chat-hf Base Model 93.80 95.82 95.25 95.87 0.00 96.21 94.56 96.04
-Golden Doc 96.33 86.80 96.28 87.78 97.11 89.42 89.69 79.29
-Random Doc 96.45 84.54 96.06 85.58 96.13 89.00 89.22 78.60
-IRM 95.78 83.33 95.77 85.09 97.06 91.67 92.96 80.47
-Llama-3-8B-Instruct Base Model 97.53 91.30 97.59 91.30 97.90 89.22 98.15 90.08
-Golden Doc 96.78 85.82 96.36 86.48 94.86 85.89 93.03 76.31
-Random Doc 96.46 83.63 96.15 84.50 94.71 84.83 96.08 81.97
-IRM 95.84 80.68 95.55 82.33 98.63 93.16 89.31 76.55
-Qwen1.5-7B-Chat Base Model 98.00 93.66 98.50 94.03 97.64 92.34 98.00 92.80
-Golden Doc 96.33 83.87 95.94 84.39 94.20 84.82 92.90 76.17
-Random Doc 95.55 80.39 95.13 80.39 93.68 82.59 95.15 83.18
-IRM 95.07 78.05 93.99 77.91 98.41 91.10 89.87 72.35
-Qwen2.5-7B-Instruct Base Model 98.52 94.54 98.98 95.42 97.15 95.54 98.62 93.80
-Golden Doc 96.25 83.91 96.00 84.83 95.33 87.23 96.24 82.33
-Random Doc 95.92 80.28 95.57 80.46 95.28 84.24 96.51 82.87
-IRM 96.04 80.20 95.31 80.68 97.75 91.73 93.78 78.70
-Table 7: Comparison of answer confidence scores for different models after robust training strategies on HotpotQA, NQ, WebQ and TriviaQA datasets.
-Model RAG Scenario HotpotQA NQ WebQuestions TriviaQA
-EM F1 EM F1 EM F1 EM F1
-Llama-2-7b-chat-hf Base Model 3.30 12.34 1.21 10.61 0.00 13.08 4.32 20.27
-RALM 26.21 36.42 27.90 38.41 30.22 46.22 55.36 63.44
-RetRobust 31.29 43.65 30.72 43.69 30.94 48.52 59.56 68.58
-Top-1 Doc 31.76 43.95 31.31 44.02 30.94 49.73 59.99 68.67
-Golden Doc 30.67 42.78 30.59 42.52 30.22 46.58 58.59 67.25
-Random Doc 30.94 43.11 31.08 43.92 30.58 48.58 60.42 69.16
-Irrelevant Doc 31.01 42.98 31.10 43.41 29.14 46.70 59.25 67.92
-IRM 34.38 47.11 30.79 43.06 31.29 49.51 59.57 68.26
-Llama-3-8B-Instruct Base Model 23.31 32.60 30.04 41.59 26.98 43.25 58.80 66.45
-RALM 27.64 38.10 30.34 41.25 30.22 44.94 58.72 66.90
-RetRobust 36.06 48.99 34.05 47.10 31.65 49.53 61.62 70.99
-Top-1 Doc 36.72 49.30 34.32 47.12 30.22 49.96 62.16 71.28
-Golden Doc 35.52 48.31 33.78 46.45 29.86 47.26 60.72 69.89
-Random Doc 35.98 49.05 34.13 47.30 28.70 47.94 61.94 71.16
-Irrelevant Doc 35.31 47.92 33.80 47.03 31.65 50.63 61.37 70.77
-IRM 35.19 48.08 34.29 47.15 30.94 49.09 61.68 70.80
-Qwen1.5-7B-Chat Base Model 24.15 34.06 24.31 35.27 23.02 38.65 48.43 57.65
-RALM 24.44 34.23 25.00 35.02 23.74 37.56 49.97 58.42
-RetRobust 29.05 40.99 25.84 37.67 27.34 43.27 52.97 61.93
-Top-1 Doc 29.68 41.57 25.21 37.12 25.18 42.39 52.27 61.90
-Golden Doc 28.66 40.55 25.43 36.84 24.46 41.78 51.26 60.34
-Random Doc 28.83 40.66 25.88 38.04 26.62 43.43 52.47 61.78
-Irrelevant Doc 28.12 39.87 24.67 36.46 25.54 42.16 51.45 60.55
-IRM 26.95 38.67 25.98 37.63 25.54 41.77 52.02 61.37
-Qwen2.5-7B-Instruct Base Model 25.10 35.02 26.97 38.15 25.90 42.46 53.86 62.73
-RALM 26.37 36.30 26.98 37.68 26.98 42.99 54.43 62.86
-RetRobust 30.47 42.26 28.58 41.25 28.78 45.61 58.52 67.21
-Top-1 Doc 30.24 42.61 28.85 41.41 28.78 46.31 58.74 67.34
-Golden Doc 30.28 41.84 28.93 41.10 28.78 46.68 57.64 66.39
-Random Doc 30.25 42.22 28.73 41.19 28.42 44.72 58.40 67.38
-Irrelevant Doc 29.76 41.67 28.40 41.00 27.70 43.90 58.40 67.38
-IRM 29.58 41.73 28.77 41.47 25.54 42.84 58.28 67.17
-Table 8: Comparison of different document selection strategies on three datasets (NQ, WebQuestions, TriviaQA) after training on HotpotQA.
-The gray cells indicate results evaluated on HotpotQA, while the orange cells indicate results evaluated on other datasets to assess generalization ability.
+6. Conclusion
+Our study reveals that as language model capacity increases, the benefits of complex robust RAG training strategies diminish significantly.
+We observe a counter-intuitive phenomenon: as model capability increases, the marginal robustness benefit of complex training diminishes sharply, and random document training often matches or exceeds the performance of complex methods.
+This effect generalizes across temporal reasoning, legal knowledge, and scenarios with conflicting evidence.
+Our analysis shows that stronger models possess inherent capabilities—better confidence calibration, cross-dataset generalization, and effective attention patterns—that reduce their reliance on complex training.
+These findings enable simplified RAG system design through streamlined retrieval mechanisms and open the door to more scalable systems that can effectively learn from noisy retrieval results without elaborate document filtering.
+Acknowledgement
+This work was supported by the Strategic Priority Research Program of the CAS under Grants No.XDB0680302, the National Natural Science Foundation of China (NSFC) under Grants No.
+62276248, the Key Research and Development Program of Xinjiang Uyghur Autonomous Region Grant No.
+2024B03026, the Beijing Nova Program under Grants No.
+20250484765, and the Youth Innovation Promotion Association CAS under Grants No.
+2023111.